### v1

# Grid-based Multi-Robot Coverage Path Planning

Abstract
Coverage Path Planning (CPP) addresses the challenge of planning paths to completely cover a designated workspace and is crucial for various robotic applications, including indoor cleaning, inspection, and outdoor tasks such as automated harvesting and environmental monitoring. Multi-Robot Coverage Path Planning (MCPP) extends this by coordinating multiple robots to enhance coverage efficiency and system robustness. This paper focuses on the grid-based MCPP problem, where the workspace is represented as a four-way connected 2D grid graph. Unlike traditional approaches that rely on the k-Traveling Salesman Problem (k-TSP) or Spanning Tree Coverage (STC), we reformulate MCPP directly on the edge-weighted graph to be covered and incorporate inter-robot conflicts and turning costs, reflecting more practical considerations in real-world applications. We propose several key improvements over our previous work: a reformulated Extended-STC (ESTC) with local optimizations, a refined local search framework (LS-MCPP), and an adaptive multi-goal planning approach combined with Priority-Based Search (PBS) for conflict resolution. Our proposed pipeline scales efficiently to large grid graphs (up to $256\times 256$) and up to 100 robots, demonstrating superior solution quality and success rates within about 18 minutes. We also validate our approach through a multi-robot system deployment, showcasing its practical effectiveness in real-world scenarios.
Index Terms:
Multi-Robot System, Automated Planning.
I Introduction
Coverage Path Planning (CPP) addresses the problem of determining a path that fully covers a designated workspace [galceran2013survey]. This problem has been extensively studied for a variety of robotic applications, ranging from indoor tasks like vacuum cleaning [wang2024apf] and inspection [almadhoun2018coverage], to outdoor activities such as automated harvesting [chen2024optimizing], planetary exploration [santra2024risk], and environmental monitoring [sudha2024coverage]. Multi-Robot Coverage Path Planning (MCPP) extends traditional CPP by coordinating the paths of multiple robots to collectively cover a workspace, thereby enhancing both the efficiency of the coverage task [tang2022learning] and the robustness of the overall system [sun2021ft]. Developing efficient and robust MCPP algorithms is crucial for the widespread deployment of multi-robot systems across various real-world applications. In this paper, we focus on a specific MCPP formulation known as grid-based MCPP, where the workspace is represented as a four-way connected 2D grid graph. The objective is for a team of robots to jointly cover all vertices of the graph, a problem analogous to the NP-hard k-Traveling Salesman Problem (k-TSP)[francca1995m]. However, solving k-TSP in the context of grid-based MCPP is challenging due to the large number of robots and the extensive graph size, which often comprises tens of thousands of vertices. Consequently, existing grid-based MCPP methods leverage the grid structure to find solutions that are constrained to the four-way connected grid space. Specifically, Spanning Tree Coverage (STC)[gabriely2001spanning] is used to generate circumnavigating coverage paths for individual robots, which sequentially follow one side of the spanning edges on a hypergraph (also known as the terrain graph) of the decomposed grid. The MCPP is then solved either by segmenting the STC path and distributing the segments among the robots[hazon2005redundancy, tang2021mstc], or by identifying the covering subtrees and generating separate STC paths for each robot [zheng2007robot, zheng2010multirobot]. For simplicity, the prefix ”grid-based” is omitted throughout the paper when referring to MCPP unless specified otherwise.
Figure 1: Simulation of 100 robots cooperatively performing a coverage task on the $256\times 256$ grid map NewYork1 [sturtevant2012benchmarks] without conflicts using our proposed planning pipeline, which reduces task completion time (i.e., makespan) by 42% compared to the baseline and resolves 556 conflicts within 18 minutes. Left: Task progress at one-third completion. Right: Task is completed.
We reformulate MCPP directly on the edge-weighted graph to be covered, explicitly considering inter-robot conflicts (i.e., collisions) to enable broader applicability in real-world robotics scenarios. The edge weights can represent various meaningful metrics, such as traversal costs on uneven terrains [qiu2024terrain] or the importance of task-specific edges that may require additional traversal time [freda20193d]. Additionally, we account for turning costs when evaluating path costs on the graph —– a critical factor in real-world coverage applications [vandermeulen2019turn, lu2022tmstc, ramesh2022optimal], as it demands additional execution time for robots. This problem setup is significantly more challenging than in the existing literature, where traversal costs are defined on hypergraph edges without considering turning costs and inter-robot conflicts. To address these challenges, we introduce several key improvements over our previous work [tang2024large], which initially computes a set of covering subgraphs for the target graph and subsequently generates a single-robot circumnavigating path for each subgraph, ensuring that each robot efficiently covers its assigned region. First, we describe the multi-robot system and formulate the problem in Sec.III. Next, we enhance the Extended-STC (ESTC) method by encoding cost differences between circumnavigating paths based on whether each hypergraph edge is used, and introduce two local optimizations (Sec.IV). With ESTC, we enhance our local search framework (LS-MCPP) to iteratively improve an initial solution. This refinement incorporates additional vertex-wise operators that facilitate better exploration of the solution space while initially ignoring inter-robot conflicts. (Sec.V). Finally, we propose an adaptive approach as a low-level multi-goal planner for each robot, combined with Priority-Based Search (PBS)[ma2019searching] as a high-level planner to resolve pairwise conflicts, together efficiently deconflicting the LS-MCPP solution (Sec.VI). Our pipeline efficiently scales to instances involving up to $256\times 256$ grid graphs and 100 robots, achieving superior solution quality and success rates compared to existing approaches within a runtime of approximately 15 minutes (Sec.VII). Moreover, we deploy a multi-robot system to demonstrate the effectiveness of our pipeline in real-world robotics applications (Sec. VII-E).
II Related Work
In this section, we survey related work on MCPP and MAPF.
II-A Multi-Robot Coverage Path Planning (MCPP)
MCPP focuses on finding a set of paths for a team of robots to achieve complete coverage of a given workspace. Following the taxonomy of single-robot CPP methods [Tomaszewski-2020-125840], we review three types of MCPP methods according to their levels of decomposition on the target workspace.
Grid-Based Methods: Grid-based methods represent the workspace as a grid graph [kapoutsis2017darp, azpurua2018multi, li2023sp2e], allowing the application of various graph algorithms. One common approach involves decomposing the graph for the robots [zheng2007robot, zheng2010multirobot, tang2023mixed] and using the Spanning Tree Coverage (STC) algorithm [gabriely2001spanning] to generate single-robot coverage paths. As mentioned in Sec. I, these methods operate on the hypergraph of the workspace, where each hypergraph vertex is decomposed into four smaller vertices. Rather than segmenting the graph directly, they decompose the hypergraph because STC relies on a spanning tree of the hypergraph to generate the single-robot coverage paths. Our previous work [tang2024large] proposed an alternative approach that operates directly on the graph to be covered, utilizing an extended version of STC to generate single-robot coverage paths on any subgraph. Unlike the hierarchical approach mentioned above, another type of approach [hazon2005redundancy, tang2021mstc] first generates a single-robot coverage path and then divides it into multiple segments, which are subsequently distributed among the robots. A significant assumption in these grid-based methods is that the hypergraph is vertex-weighted, with the weight of each hypergraph vertex evenly distributed across the four decomposed vertices. The cost of each coverage path is evaluated by summing the weights of all the decomposed vertices along the path, which may not accurately reflect the actual traversal costs for the robots, as these weights heavily constrain them. A more general and effective practice is to directly define the traversal cost based on the weights of the edges traversed by the robots, as in our formulation in our formulation in Sec. III-A), which, in the case of grid-based MCPP, corresponds to the edges of the graph to be covered.
Cellular Decomposition Methods: Cellular decomposition methods [rekleitis2008efficient, xu2014efficient, karapetyan2017efficient] divide the workspace into sub-regions by detecting geometric critical points, using techniques such as trapezoidal decomposition [latombe1991exact], boustrophedon decomposition [choset2000coverage], and Morse decomposition [acar2002morse]. After the division, a back-and-forth patterned path is generated to cover each sub-region, with possible optimization of the path alignment orientation [bochkarev2016minimizing]. While efficient, these methods are less suited for obstacle-rich or non-rectil inear workspaces due to their reliance on geometric partitioning. Additionally, they assume uniform traversal costs across the workspace, making them inapplicable in scenarios where robots require nonuniform time to move or perform specific coverage tasks.
Global Methods: Global methods are primarily studied and applied in additive manufacturing [yang2002equidistant, ren2009combined, gibson2021additive] (also known as layered fabrication), where a 3D object is printed layer by layer. Each layer represents a coverage problem, requiring the planning of a ”space-filling” curve for the nozzle to cover the 2D plane. Due to the path curvature requirements in these applications, global methods do not explicitly decompose the workspace; instead, they directly generate paths to cover the workspace, minimizing path curvature and discontinuity. Recent works have extended the connected Fermat spiral [zhao2016connected] to reduce energy consumption for CPP on undulating terrains [wu2019energy] and to minimize the makespan for general MCPP tasks [tang2024multi]. The coverage quality (e.g., coverage and overlapping ratios) of global methods depends on the nature of the generated space-filling curve, often requiring additional trajectory optimization as a post-processing step. A further challenge for global MCPP methods is task distribution among robots for makespan minimization, as the space-filling curve typically forms a closed loop with two endpoints.
II-B Multi-agent Path Finding (MAPF)
MAPF seeks to generate a set of plans for multiple agents to move from start to goal locations without conflicts (i.e., collisions) with other agents. For detailed surveys on MAPF, we refer readers to [stern2019multi, stern2019multi2, ma2022graph]. Here, we review the multi-goal variants of MAPF most relevant to this paper.
A recent study formulates an ordered multi-goal MAPF problem (OMG-MAPF) [mouratidis2024fools] to find plans for multiple agents so that each agent visits an ordered set of goal locations without conflicts. The order of goal locations for each agent is known as the total order. A related problem, MAPF-PC [zhang2022multi], considers precedence constraints on pairs of goal locations for each agent, meaning one goal location must be visited before the other. This precedence constraint is referred to as the partial order, and partial orders may only apply to specific pairs of agents in MAPF-PC. A set of partial orders can be converted into a total order via topological sorting, and a total order can also be extracted from a set of partial orders. For both OMG-MAPF and MAPF-PC, it is assumed that all actions have unit costs and that agents operate on a clock discretized into unit time steps. Thus, these methods are not directly applicable to graphs with nonuniform edge weights.
Another line of work on multi-goal variants, called MG-MAPF [surynek2021multi, ren2021ms, ren2022conflict, tang2024mgcbs], considers an unordered set of goal locations for each agent, with the additional objective of determining the optimal global ordering of goal locations (i.e., optimal goal sequencing) to minimize an objective function. Although one might formulate MCPP (as defined in Problem III) as an MG-MAPF problem, co-optimizing the goal sequencing for the vertices along each coverage path might not yield a better sequence than the original coverage path within a limited planning time, given the combinatorial complexity of co-optimization. Intuitively, as the robot must start and return to its root vertex in MCPP, most conflicts can be resolved by short-duration wait actions without the need to re-sequence the vertices along a coverage path.
The Multi-agent Pickup and Delivery (MAPD) problem [ma2017lifelong, ma2019lifelong, li2021lifelong, xu2022multi] can also be considered a multi-goal MAPF variant, where a system continuously assigns new tasks to agents. However, agents are unaware of each task’s pickup and delivery locations before the task is assigned, making it more difficult to address the goal sequencing problem.
III Problem Formulation
In this section, we first describe the multi-robot system with several key assumptions. Upon this, we address the MCPP problem through a progressive approach by solving a relaxed version of MCPP that disregards inter-robot conflicts and then resolving the conflicts among the generated coverage paths.
III-A Description of the Multi-Robot System
We consider a set $I=\{1,2,...,k\}$ of $k$ homogeneous and holonomic robots that operate in a four-way connected undirected 2d grid graph $G=(V,E)$, where $V$ and $E$ are the vertex set and edge set, respectively. Each vertex $v\in V$ is connected by edges to its top, bottom, left, and right neighbors if they exist. Each edge $e\in E$ is associated with a non-negative weight $w_{e}$, and we simply set $w_{e}=1$ for every $e\in E$ when considering an unweighted graph $G$.
We represent the robot state of arriving vertex $v\in V$ with a heading $\psi$ at time $t$ as $\mathbf{x}=(v,t,\psi)$. As graph $G$ is a four-way connected grid graph, we restrict the heading $\psi\in\{\text{E},\text{N},\text{W},\text{S}\}$ to East (E), North (N), West (W), and South (S). We assume the system has perfect knowledge about the robot states and put our focus of this paper on the time-optimal planning problems given in the next two subsections. Each robot $i$ is initialized in the state $(r_{i},0,\text{N})$, where $r_{i}\in V$ is called the root vertex given as an input. The robots can take the following three types of actions to transit from a state $\mathbf{x}=(v,t,\psi)$ to another state $\mathbf{x}^{\prime}=(v^{\prime},t^{\prime},\psi^{\prime})$:
  • Wait action: a robot waits at vertex $v$ for a time period $t^{\prime}-t>0$ where $v=v^{\prime}$ and $\psi=\psi^{\prime}$.
  • Turn action: a robot turns at vertex $v=v^{\prime}$ from $\psi$ to $\psi^{\prime}\neq\psi$ at the time cost of $t^{\prime}-t=c_{t}(\psi,\psi^{\prime})>0$, where $c_{t}$ is a function defined as $c_{t}(\psi,\psi^{\prime})=C\cdot\frac{|\psi^{\prime}-\psi|}{90^{\circ}}$. The constant $C$ is specified depending on the robots.
  • Move action: a robot moves from vertex $v$ to vertex $v^{\prime}\neq v$ via $e=(v,v^{\prime})\in E$ at the time cost of $t^{\prime}-t=w_{e}>0$.
In the rest of the paper, we will use tuple $(G,I,R)$ to denote the above Multi-Robot System (MRS) with a set $I$ of $k$ robots that live on the graph $G=(V,E)$ and are initialized in the set $R=\{r_{i}\}_{i=1}^{k}\subseteq V$ of root vertices. We aim to generate a set of plans for robots in the MRS to solve the problems in Sec. III-B, where a plan for any robot is defined as follows:
Definition 1 (Plan).
A plan $P=(\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{|P|})$ for the robot $i\in I$ is a finite ordered sequence of states such that $\mathbf{x}_{1}=(r_{i},0,\text{N})$, $\mathbf{x}_{|P|}=(r_{i},t_{\text{final}},\cdot)$, and every pair of adjacent states are in transit through wait, turn, or move actions.
The above definition follows the cover with return coverage problem setting [zheng2010multirobot], where the robots must start from and return to their corresponding initial vertices. The time cost $c(P)=t_{\text{final}}$ of plan $P$ is obtained from its last state $\mathbf{x}_{|P|}$. For convenience, we use $P[j]$ to represent the $j$-th state $\mathbf{x}_{i}$ of any plan $P$; when $j$ is negative, $P[j]$ represents the reversed $|j|$-th state $\mathbf{x}_{|P|+j+1}$ of $P$. We further denote the set of unique vertices visited by $P$ as $V(P)\subseteq V$, and the list of (potentially duplicated) edges used by $P$ as $E(P)$. To take the inter-robot conflict avoidance into consideration of planning, we formally define the conflict between any pair of plans as follows:
Definition 2 (Conflict).
Given any two plans $P_{a}$ and $P_{b}$, a conflict on vertex $v\in V$ occurs if there exists two states $P_{a}[i]=(v,t_{i},\cdot),P_{a}[i+1]=(\cdot,t^{\prime}_{i},\cdot)$ and two states $P_{b}[j]=(v,t_{j},\cdot),P_{b}[j+1]=(\cdot,t^{\prime}_{j},\cdot)$, such that the two time intervals $[t_{i},t^{\prime}_{i})$ and $[t_{j},t^{\prime}_{j})$ intersects with each other.
Our definition of conflict is more conservative than that in most MAPF algorithms. Specifically, we assume that a vertex is blocked not only when it is occupied by a robot, as is typical in the MAPF literature [sharon2015conflict], but also when a robot is moving out of it. By checking for time intervals during which other robots occupy a vertex, we can implicitly identify all conflict types outlined in the literature [stern2019multi], including vertex conflicts, edge conflicts, following conflicts, and swapping conflicts. These conflicts are typically defined for unweighted graphs, where time is discretized into uniform intervals for the MRS, allowing their actions to be aligned step by step. In addition, we adopt the conservative assumption of wellformness as defined in Definition 3 following the approach of [ma2019searching], which ensures that PBS-like algorithms are complete (i.e., always produce a valid solution) for various MAPF variants.
Definition 3 (Wellformness).
In a wellformed MRS $(G,I,R)$, each robot $i\in I$ can wait indefinitely at its root vertex $r_{i}\in V$ without blocking any other robot.
III-B Grid-based Multi-Robot Coverage Path Planning
For an MRS $(G,I,R)$ described above, we formally define the MCPP problem for the MRS as below:
Problem 1 (MCPP).
Find a set $\mathcal{P}=\{P_{i}\}_{i=1}^{k}$ of plans for the set $I$ of robots such that $\bigcup_{i=1}^{k}V(P_{i})=V$ and there is no conflict between any pair of plans in $\mathcal{P}$.
Figure 2: The proposed algorithmic pipeline for MCPP (Problem 1) operates by first solving Relaxed-MCPP (Problem 2) to obtain a set of coverage paths, with which as input to further solve Deconflicted-MCPP (Problem 3) to obtain a set of plans, resulting the final set of action sequences for the robots.
We use makespan metric to evaluate the quality of an MCPP solution, denoted as $\tau(\mathcal{P})=\max\{c(P_{1}),c(P_{2}),...,c(P_{k})\}$. We emphasize that our formulation of the MCPP problem, as defined in Problem 1, differs from existing approaches in three key aspects: (1) it directly uses any edge-weighted graph as input, rather than the hypergraph of the graph to be covered; (2) it incorporates turning costs into the overall plan cost, not just traversal costs; and (3) it accounts for inter-robot conflicts. These distinctions significantly increase the difficulty of finding an optimal conflict-free solution, as we will demonstrate in Sec. III-C, where we show that optimally solving its two sub-problems are already NP-hard. To enhance the computational tractability of optimally solving Problem 1, we decompose it into two sub-problems. The first sub-problem addresses a relaxed version of MCPP by finding a set of paths that cover all vertices in $G$. These paths are specifically defined for the robots in the MRS $(G,I,R)$ as follows:
Definition 4 (Path).
A path $\pi=(v_{1},v_{2},...,v_{|\pi|})$ for the robot $i\in I$ is a finite ordered sequence of vertices such that $v_{1}=v_{|\pi|}=r_{i}$ and $(v_{j-1},v_{j})\in E$ for each $j=2,3,...,|\pi|$.
Similarly to the notation for a plan, we use $V(\pi)$, $E(\pi)$, and $\pi[j]$ to denote the set of unique vertices in $\pi$, the list of edges used by $\pi$, and the $j$-th vertex in $\pi$, respectively. The relaxed version of Problem 1 for MRS $(G,I,R)$, which corresponds to the classic MCPP problem without considering inter-robot conflicts as discussed in [tang2024large], is formally defined as follows:
Problem 2 (Relaxed-MCPP).
Find a set $\Pi=\{\pi_{i}\}_{i=1}^{k}$ of paths such that $\bigcup_{i=1}^{k}V(\pi_{i})=V$.
After finding a set $\Pi$ of paths for Problem 2, we use it as input to solve the second sub-problem, which aims to construct a set $\mathcal{P}$ of conflict-free plans for Problem 1. Formally, the second sub-problem for MRS $(G,I,R)$ is defined as below:
Problem 3 (Deconflicted-MCPP).
Given a Problem 2 solution $\Pi=\{\pi_{i}\}_{i=1}^{k}$, find a set $\mathcal{P}=\{P_{i}\}_{i=1}^{k}$ of plans such that (1) there is no conflict between any pair of plans in $\mathcal{P}$; (2) for each $i\in I$, there exists a subsequence $(P_{i}[l_{1}],P_{i}[l_{2}],...,P_{i}[l_{|\pi_{i}|}])$ of $P_{i}$ where each $P_{i}[l_{j}]=(\pi_{i}[j],\cdot,\cdot)$ and each $l_{j}<l_{j+1}$.
In Problem 2, the paths from set $\Pi$ essentially disregard inter-robot conflicts and assume instantaneous heading changes between vertices, without considering intermediate turning actions for the robots. Consequently, for each robot $i\in I$, we can compute a cost $c(\pi_{i})$ as a lower bound of the true plan cost $c(P_{i})$ that accounts for conflicts as follows:
$c(\pi_{i})=\sum_{e\in E(\pi_{i})}w_{e}+\sum_{j=2}^{|\pi_{i}|}c_{t}(\psi_{j-1},\psi_{j})\leq c(P_{i})$ (1)
where each heading $\psi_{j}\in\{E,N,W,S\}$ is determined by the relative position between $\pi_{i}[j-1]$ and $\pi_{i}[j]$. As a lower bound of $c(P_{i})$, $c(\pi_{i})$ can guide the solution of Problem 2 by optimizing the makespan $\tau(\Pi)=\max\{c(\pi_{1}),c(\pi_{2}),...,c(\pi_{k})\}$. Ultimately, this approach aims to construct a makespan-minimized plan for Problem 1 by solving Problem 3. As shown in Fig. 2, in this paper, we propose to solve MCPP by sequentially solving Relaxed-MCPP and Deconflicted-MCPP.
III-C Theoretical Analysis
We show that optimally solving each sub-problem of MCPP is NP-hard, necessitating our time-efficient suboptimal algorithms in Sec. V and VI to solve them separately.
Theorem 1.
It is NP-hard to find an optimal Relaxed-MCPP solution with a minimized makespan.
Proof 1.
Inspired by [zheng2010multirobot], we prove Theorem 1 by reducing the strongly NP-hard 3-Partition problem (3-PP) [gary1979computers] to the decision version of Relaxed-MCPP, that is, determining whether Relaxed-MCPP has a solution with its makespan larger than a given positive number. Specifically, we consider a positive integer $B$ and a set $S=\{a_{i}\}_{i=1}^{3k}$ of positive integers with each $a_{i}\in(\frac{B}{4},\frac{B}{2})$ and $\sum_{i=1}^{3k}a_{i}=kB$, the objective of the 3-PP is to determine whether we can evenly partition $S$ into $k$ subsets $S_{1},S_{2},...,S_{k}$ of indices. We now construct a Relaxed-MCPP instance with the gird graph as shown in Fig. 3, where there is a horizontal “corridor” with $4k$ vertices $r_{n},r_{n-1},...,r_{1},v_{1},v_{2},...,v_{3k}$ at the center and $3k$ “tunnel” vertices $u_{1},u_{3},...,u_{3k}$ each connecting $v_{i}$. If index $i$ is odd, $u_{i}$ is on the top of $v_{i}$ otherwise on the bottom. The root vertex of robots $1,2,...,k$ are set to $r_{1},r_{2},...,r_{k}$, respectively. Each edge connecting corridor vertices have a weight of 1, whereas each tunnel vertex connecting the corridor vertex $v_{i}$ have a weight of $4ka_{i}$. As the instance construction can be completed in polynomial time, we now prove the claim that the makespan of the Relaxed-MCPP instance is at most $28kB+8k-$ 2 iff $S$ can be evenly partitioned into $k$ subsets:
$\bullet$ “If” direction: For each subset $S_{i}$ of the even partition, we know that $\sum_{i\in S_{j}}a_{i}=B$. Let the robot $j$ cover each tunnel vertex $u_{i}$ where $i\in S_{j}$, then its time cost is $\sum_{i\in S_{j}}4ka_{i}\times 2=8kB$ to cover the tunnel vertices and at most $(4k-1)\times 2$ to traverse through the corridor (from $r_{k}$ to $v_{3k}$). Combining both we have the time cost of each robot is at most $28kB+8k-$ 2 and thus the makespan is at most $28kB+8k-$ 2.
Figure 3: A Relaxed-MCPP instance reducible from the 3-Partition problem.
$\bullet$ “Only if” direction: We define $S_{j}$ to be the set of indices such that each $u_{i}$ for $i\in S_{j}$ is first covered by robot $j$, resulting in a partition of $k$ subsets $S_{1},S_{2},...,S_{k}$. As each $a_{i}\in(\frac{B}{4},\frac{B}{2})$ and $|S|=3k$, we know that each robot $j$ covers exactly three tunnel vertices and thus its time cost is at least $\sum_{i\in S_{j}}(4ka_{i}\times 2)+3\times 2=8k\sum_{i\in S_{j}}a_{i}+6$. Since the makespan is at most $28kB+8k-$ 2, we know the time cost of each robot $j$ is at most $28kB+8k-$ 2. As $a_{i}$, $B$ and $k$ are all positive integers, we have $\sum_{i\in S_{j}}a_{i}\leq B+1-\frac{1}{k}\leq B$ by chaining the above two inequalities. Since $\sum_{j=1}^{k}\sum_{i\in S_{j}}a_{i}=\sum_{i=1}^{3k}a_{i}=kB$, we have $\sum_{i\in S_{j}}a_{i}=B$ must hold to satisfy $\sum_{i\in S_{j}}a_{i}\leq B$, and thus $S_{1},S_{2},...,S_{k}$ is an even partition.
Theorem 2.
It is NP-hard to find an optimal Deconflicted-MCPP solution with a minimized makespan.
Proof 2.
The OMG-MAPF problem (see Sec. II-B) was proved to be NP-hard with a makespan minimization objective [mouratidis2024fools], which can be viewed as a special case of Deconflicted-MCPP operating on unweighted graphs (i.e., $w_{e}=1$ for each $e\in E$). Therefore, it is at least NP-hard to find an optimal Deconflicted-MCPP solution with a minimized makespan.
IV Single-Robot Coverage Path Planning
In this section, we address the single-robot CPP problem, which is a specific case of MCPP (Problem LABEL:problem:_mcpp) when $k=|I|=1$ in instance $(G,I,R)$. Solving CPP efficiently is fundamental to our local search framework for MCPP in Sec. V. We propose Extended-STC (ESTC) to solve CPP, which builds on the STC algorithm for offline CPP and Full-STC [gabriely2002spiral] for online CPP. Both STC and Full-STC treat $G$ as a vertex-weighted graph to evaluate the traversal cost, where the weight of each vertex $v\in V$ depends on a hypervertex that can be decomposed into $v$ to reserve the solution optimality. However, this special treatment does not fully account for the actual traversal time cost incurred by a robot, which should be more naturally captured by the edges the robots traversed.
IV-A Extended Spanning Tree Coverage (ESTC)
ESTC requires constructing an auxiliary hypergraph $H=(V_{h},E_{h})$ from graph $G$ by contracting neighboring vertices within each grid cell into a hypervertex and connecting them with hyperedges between adjacent hypervertices. As shown in Fig. 4, the cells, represented by the light grey squares, are formed by contracting black vertices from $V$ into hypervertices. The graph $G$ is contracted left to right and bottom to top by default during the hypergraph construction. For clarity, we let $\delta_{v}\in V_{h}$ denote the hypervertex contracted from $v\in V$ and $\varepsilon=(\delta_{u},\delta_{v})\in E_{h}$ denote the hyperedge connecting $\delta_{u}$ and $\delta_{v}$. A hypervertex $\delta$ is termed incomplete if it is formed from fewer than four vertices; otherwise, it is complete. A special case arises when only two diagonal vertices exist within a grid cell —– each diagonal vertex forms its own nonadjacent hypervertex since they are not directly connected. To define the hyperedge weights for $H$, we categorize each hyperedge $\varepsilon=(\delta_{u},\delta_{v})\in E_{h}$ based on the number of crossing edges between the corresponding vertices within the cells of $\delta_{u},\delta_{v}$. As illustrated in Fig. 4, a crossing edge is an edge $(u,v)\in E$ where $u$ and $v$ are contracted into $\delta_{u}$ and $\delta_{v}$, respectively. The hyperedge weight $w_{\varepsilon}$ is then defined as follows:
$w_{\varepsilon}=\begin{cases}w_{(u_{1},v_{1})}+x_{(u_{2},v_{2})}-w_{(u_{1},u_{2})}-w_{(v_{1},v_{2})},&x_{\varepsilon}=2\\ 2\cdot w_{(u_{1},v_{1})},&x_{\varepsilon}=1\end{cases}$ (2)
where $x_{\varepsilon}$ is the number of crossing edges in $\varepsilon$. When $x_{\varepsilon}=0$, we know that $(\delta_{u},\delta_{v})\notin E_{h}$ and thus do not need to define its edge weight. Intuitively speaking, $w_{\varepsilon}$ implicitly encodes the cost difference between two circumnavigating coverage paths without explicitly generating them — one includes the crossing edges in $\varepsilon$ and the other does not. By Theorem 4, we demonstrate that Eqn. (2) ensures that ESTC (Alg. 1) always finds an optimal circumnavigating coverage path.
Figure 4: Illustration of crossing edges in hyperedge $\varepsilon=(\delta_{u},\delta_{v})\in E_{h}$.
We describe the pseudocode of ESTC in Alg. 1. After the hypergraph construction [Line 1], ESTC reduces the CPP problem to finding the Minimum Spanning Tree (MST) of $H$. Specifically, we use Kruskal’s algorithm to find the MST [Line 1] as the hyperedge weight could be negative. Then akin to STC, ESTC computes the coverage path by circumnavigating each MST edge via a Depth-First-Search [Line 1]. While existing STC-based CPP and MCPP methods [zheng2010multirobot, tang2023mixed] assume all hypervertices are complete, ESTC integrates routing rules from Full-STC to handle situations where hypervertices are incomplete, which redirects the path to the other side of a spanning edge when it is infeasible to move along one side due to the nonadjacency of vertices within the incomplete hypervertices. This makes ESTC a general algorithm to be applied on an arbitrary grid graph without assuming that all the hypervertices are complete, which is essential to the local search framework for MCPP in Sec. V.
Input: grid graph $G=(V,E)$
1 construct hypergraph $H=(V_{h},E_{h})$ from $G$
2 $M\leftarrow$ find MST of $H$ via Kruskal’s algorithm
3 $\pi\leftarrow$ compute the circumnavigating path of $M$ with the Full-STC routing rules via Depth-First-Search
4 return the coverage path $\pi$
Algorithm 1 Extended-STC (ESTC)
IV-B Theoretical Analysis
This subsection shows that ESTC achieves complete coverage in Theorem 3 and is the optimal circumnavigating coverage path in Theorem 4. We demonstrate that ESTC is bounded suboptimal for CPP (including non-circumnavigating paths) in Theorem 5. We further analyze the time complexity of ESTC in Theorem 6, highlighting its practical time efficiency.
Theorem 3.
ESTC achieves complete coverage on a connected graph $G$.
Proof 3.
The proof is analogous to the argument used in the proof of Lemma 3.1 in [gabriely2002spiral] given that $G$ is connected.
Theorem 4.
The cost of the circumnavigating path from a minimum spanning tree of the hypergraph is always smaller than that from any non-minimum spanning tree.
Proof 4.
For hypergraph $H=(V_{h},E_{h})$, consider a minimum spanning tree $M=(V_{h},F\cup\{\varepsilon_{m}\})$ of $H$ and a non-minimal spanning tree $N=(V_{h},F\cup\{\varepsilon_{n}\})$ of $H$. Denote $c(\pi_{M})$ and $c(\pi_{N})$ as the costs of the two circumnavigating paths resulting from $M$ and $N$, respectively. As illustrated in Fig. 4, we know that $E(\pi_{M})\setminus E(\pi_{N})=\{(u_{1},v_{1}),(u_{2},v_{2})\}$ (blue solid lines) and $E(\pi_{N})\setminus E(\pi_{M})=\{(u_{1},u_{2}),(v_{1},v_{2})\}$ (red solid lines) if $x_{\varepsilon_{m}}=x_{\varepsilon_{n}}=2$. Therefore, we can compute the cost difference as $\Delta c=c(\pi_{M})-c(\pi_{N})$ as follows:
$\Delta c$ $=[w_{(u_{1},v_{1})}+w_{(u_{2},v_{2})}-w_{(u_{1},u_{2})}-w_{(v_{1},v_{2})}]$
$-[w_{(p_{1},q_{1})}+w_{(p_{2},q_{2})}-w_{(p_{1},q_{2})}-w_{(p_{1},q_{2})}]$
Substituting Eqn. (2) into the above equation, we have $\Delta c=w_{\varepsilon_{m}}-w_{\varepsilon_{n}}<0$ as $w_{\varepsilon_{m}}<w_{\varepsilon_{n}}$ by definition. As a result, we have $c(\pi_{M})<c(\pi_{N})$ and the same derivation applies to all other combinations of $(x_{\varepsilon_{m}},x_{\varepsilon_{n}})\in\{1,2\}\times\{1,2\}$. By replacing an edge from $M$ to an edge from an arbitrary non-minimum spanning tree $T$ one at a time, we can sequentially prove $c(\pi_{M})<c(\pi_{T})$ using the above derivation and thus conclude the proof.
Theorem 5.
ESTC is bounded suboptimal for CPP.
Proof 5.
Considering the CPP problem for the MRS $(G,I,R)$, we analyze the worst-case cost of solution $\pi$ by ESTC compared with the optimal CPP solution $\pi^{*}$. Suppose that graph $G$ is contracted into hypergraph $H=(V_{h},E_{h})$ with $n_{i}$ incomplete hypervertices. To ensure that $\pi^{*}$ is a valid complete coverage path, we know that $V(\pi)=V$ and the first and last vertices of $\pi$ are both the root vertex. It follows that one can easily derive $|E(\pi^{*})|\geq|V|$, where the equality holds iff $n_{i}=0$. Then we can obtain the optimality ratio $\rho$ for ESTC:
$\rho=\frac{c(\pi)}{c(\pi^{*})}=\frac{\sum_{e\in E_{\pi}}}{\sum_{e\in E_{\pi}^{*}}}\leq\frac{|E_{\pi}|\cdot w_{\text{max}}}{|E_{\pi}^{*}|\cdot w_{\text{min}}}\leq\frac{|E_{\pi}|\cdot w_{\text{max}}}{|V|\cdot w_{\text{min}}}$ (3)
where $w_{\text{max}}$ and $w_{\text{min}}$ are the maximal and minimal weights among all hyperedges from $E_{h}$, respectively. If $n_{i}=0$, we have $|E_{\pi}|=|V|$ following the same derivation as above for $\pi^{*}$. If $n_{i}\neq 0$, ESTC adds at most two additional edges that duplicate with existing edges compared to the case of $n_{i}=0$, to traverse each incomplete hypervertex. As a result, we have $|E(\pi)|\leq|V|+2\cdot n_{i}$, with which we can further refine Eqn. (3) as follows:
$\rho\leq(1+2\cdot\frac{n_{i}}{|V|})\cdot\frac{w_{\text{max}}}{w_{\text{min}}}$ (4)
Clearly, the right-hand side of Eqn. (4) depends only on the input $G$, implying that ESTC is bounded suboptimal.
Remark 1.
The above bound of the optimality ratio is rather loose. Since $n_{i}\in[0,|V|]$, we know that the first multiplier $(1+2\cdot\frac{n_{i}}{|V|})\in[1,3]$. In practice, when $n_{i}$ is large, the optimal solution $\pi^{*}$ must adopt rerouting rules (e.g., Full-STC) to revisit edges through incomplete hypervertices for complete coverage. This puts $|E(\pi^{*})|$ close to $|E(\pi)|$ instead of $|V|$, which narrows the bound in Eqn. (3). Specifically, ESTC is optimal ($\rho=1$) if $G$ is unweighted ($\frac{w_{\text{max}}}{w_{\text{min}}}=1$) and all hypervertices are complete ($n_{i}=0$).
Theorem 6.
The time complexity of ESTC (Alg. 1) for any grid graph $G=(V,E)$ is $O(|V|\log|V|)$.
Proof 6.
To construct hypergraph $H=(V_{h},E_{h})$ from $G=(V,E)$ as described in Sec. IV-A, we iterate $V$ to build $V_{h}$ and then iterate $V_{h}$ to build $E_{h}$. As $|V_{h}|$ is proportional to $|V|$, the time complexity of this construction is $O(|V|)$. To find the MST of $H$, ESTC uses Kruskal’s algorithm therefore having a time complexity of $O(|E|\log|E|)$. The procedure to generate the circumnavigating coverage path on the MST via Depth-First-Search entails a time complexity of $O(|V|+|E|)$. As $|E|$ is proportional to $|V|$ in the grid graph $G$, combining the above, the time complexity of ESTC is $O(|V|\log|V|)$.
IV-C Local Optimizations
We introduce two efficient local optimizations for ESTC, namely Parallel Rewiring and Turn Reduction. Parallel rewiring operates as a post-processing procedure to enhance an ESTC solution, while turn reduction introduces a secondary objective during the execution of Alg. 1.
Parallel Rewiring: As the original STC only explores the solution space of circumnavigating coverage paths, we develop two parallel rewiring procedures to refine any ESTC solution $\pi$ with non-circumnavigating local paths. While they do not theoretically improve the upper bound on the suboptimality of ESTC in Theorem 5, empirical results (see Sec. IV-A) often indicate an improvement in solution quality in practice. Type-A Parallel Rewiring searches for every subsequence in $\pi$ matching the pattern $(...,v,u,...,v,t,s,...,t,...)$ and collapses it into $(...,v,u,...,s,...t,...)$ if the resulting cost difference $w_{(u,s)}-[w_{(u,v)}+w_{(v,t)}+w_{(t,s)}]<0$. As shown in Fig. 5-(a), each collapsing operation effectively removes the intermediate vertices $v,t\in V$ without losing any vertex coverage. Type-B parallel rewiring identifies every pair of two vertically (or horizontally) parallel edges $(u_{1},u_{2}),(v_{1},v_{2})\in E(\pi)$ where $(u_{1},u_{3}),(u_{2},u_{4}),(s_{1},u_{3}),(s_{2},u_{4})\in E(\pi)$ are aligned horizontally (or vertically). Fig. 5-(b) shows that it removes the edges $(u_{1},u_{3}),(u_{2},u_{4}),(v_{1},v_{2})$ and inserts the edges $(u_{3},u_{4}),(u_{1},v_{1}),(u_{2},v_{2})$ if the cost difference after rewiring is less than 00, that is, $[w_{(u_{3},u_{4})}+w_{(u_{1},v_{1})}+w_{(u_{2},v_{2})}]-[w_{(u_{1},u_{3})}+w_{(u_{2},u_{4})}+w_{(v_{1},v_{2})}]<0$. Both types have a runtime complexity of $O(|V|^{2})$ with a naive implementation of two-layer iterations on $V$.
Turn Reduction: ESTC can be seamlessly integrated with a secondary objective of turn reduction in addition to its primary objective of path cost minimization. Specifically, we consider aligning the hyperedges when constructing the MST for ESTC. As the hypergraph $H=(V_{h},E_{h})$ is a four-way-connected 2d grid graph, we can categorize the hyperedges from $E_{h}$ by their directions as vertical and horizontal. Given a goal direction to align with, we prioritize the hyperedges traversal in Kruskal’s algorithm (line 1 of Alg. 1) by (1) their directions are identical to the goal direction and (2) the sum of their two endpoints’ degrees is smaller. The first prioritization is straightforward to align hyperedges with the goal direction as much as possible, whereas the second prioritization tries to first grow the MST into those less connected regions of $H$ for better alignment. As it only introduces two priority keys when constructing the MST, ESTC with turn reduction has the same time complexity as demonstrated in Theorem 6. Fig. 6 shows an example of applying turn reduction on an ESTC coverage path.
Figure 5: Local views of ESTC paths before and after parallel rewiring.
IV-D Integrating ESTC into Existing Grid-based MCPP Methods
While existing grid-based MCPP methods mostly focus on the special case of MCPP with strict assumptions of complete hypervertices and vertex-weighted graphs, we want to emphasize our more general formulated MCPP in Problem LABEL:problem:_mcpp is relatively unexplored in the MCPP literature, despite its significant practical applications. The proposed ESTC can be seamlessly integrated into these methods, allowing them effectively solve MCPP without the aforementioned assumptions. Considering MCPP for MRS $(G,I,R)$, we show several examples of adapting them for MCPP. Comprehensive numerical results for solving MCPP using these methods are provided in Sec. VII.
VOR: VOR is a naive baseline method introduced in [tang2024large]. It first decomposes $G$ into $k$ subgraphs $\{G_{i}\}_{i=1}^{k}$ as a Voronoi diagram [aurenhammer2000voronoi], with the set $R$ of root vertices as the pivot points and the shortest path distances on $G$ as the distance metric. Then, it generates coverage path $\pi_{i}$ on each subgraph $G_{i}$ via ESTC, constituting a solution for MCPP.
MFC: Multi-robot Forest Coverage (MFC)[zheng2007robot, zheng2010multirobot] heuristically computes a set $\{T_{i}\}_{i=1}^{k}$ of $k$ covering subtrees for the hypergraph $H$ using rooted-tree-cover[even2004min], where each hypervertex is included in at least one subtree. It then generates the coverage path on each $T_{i}$ via STC. To adapt it for MCPP, we make two modifications: (1) hypergraph $H$ is constructed as per Sec. IV-A instead of hypervertex-weighted in the original MFC and (2) STC is replaced with ESTC.
MIP: MIP [tang2023mixed] follows a similar two-level approach as MFC with one key difference. Instead of using rooted-tree-cover to find the subtree cover, it solves a mixed-integer program to find an optimal subtree cover that minimize the maximal subtree weight. By applying the same modifications to MIP as MFC, it can be used to solve MCPP.
MSTC: MSTC [tang2021mstc] is an improved version of MSTC [hazon2005redundancy], which operates differently from the above approach. MSTC first computes a single-robot coverage path via STC, then segments the path into $k$ segments each corresponding to a robot. MSTC then uses a shortest path algorithm to connect the root vertex with the cut points for each robot, forming the final coverage path for each robot. By replacing STC with ESTC, we can adapt MSTC to solve MCPP.
Figure 6: ESTC coverage path before (left) and after (right) turn reduction with the edges prioritized to align horizontally when constructing the MST.
V The LS-MCPP Framework
Figure 7: The flowchart illustrates our LS-MCPP algorithmic framework on instance floor-small. Given an initial MCPP solution $\Pi_{0}$, LS-MCPP iteratively samples a pool $O$ from three pools of dedicated operators and then samples an operator $o$ from $O$, to transform a set $\mathcal{G}$ of subgraphs from one to another. It then applies ESTC on the subgraphs to evaluate the new MCPP solution $\Pi$, which is adaptively accepted according to its makespan $\tau$. A periodic subroutine of forced deduplication is called at the end of each iteration, and the best solution $\Pi^{*}$ is recorded if $\Pi$ is better with a lower makespan.
In this section, we solve MCPP defined in Problem LABEL:problem:_mcpp leveraging ESTC as a subroutine in our proposed local search algorithmic framework, namely LS-MCPP. Local search is a heuristic optimization method that directly explores the solution neighborhood to find a good solution to the problem. As depicted in Fig. 7, given an initial solution for MCPP instance $(G,I,R)$, its corresponding set $\mathcal{G}_{0}=\{G_{i}=(V_{i},E_{i})\}_{i=1}^{k}$ of $k$ connected subgraphs are computed. LS-MCPP operates as an iterative procedure utilizing a hierarchical sampling scheme to sample an operator to progressively transform the subgraph set $\mathcal{G}$ from one to another. It then employs ESTC on each subgraph $G_{i}$ to generate its coverage path and evaluate the makespan of MCPP. In the following subsections, we first provide an overview of LS-MCPP and then detailed descriptions of its key components.
V-A Algorithm Description
Overview: Consider the MCPP instasnce $(G,I,R)$, LS-MCPP incorporates three types of operators to alter a set $\mathcal{G}=\{G_{i}\}_{i=1}^{k}$ of subgraphs. We classify each subgraph $G_{i}$ as a light subgraph if its coverage path cost $c(\pi_{i})\leq\bar{c}$ by applying ESTC on each subgraph, where $\bar{c}$ is the average cost of all the coverage paths. Conversely, if its cost is greater than $\bar{c}$, we call $G_{i}$ a heavy subgraph. LS-MCPP utilizes the grow operators on light subgraphs to assign new vertices to cover and deduplicate operators on heavy subgraphs to eliminate unnecessarily repeated coverage. In addition, it employs exchange operators to balance the costs between subgraphs with a high cost difference. The operators are introduced in Sec. V-B, whose aims are to attain a cost-equilibrium MCPP solution with a low makespan. In general, LS-MCPP adopts a two-level hierarchical sampling approach to efficiently explore the constructed neighborhood in an iterative fashion. The first layer selects an operator pool using roulette wheel selection [goldberg1989genetic] from three pools, each containing a specific type of operator. The second layer heuristically draws an operator from the pool selected in the first layer, as detailed in Sec. V-C.
Input: MCPP instance $(G,I,R)$, initial solution $\Pi_{0}$
Param: max iteration $M\in\mathbb{Z}^{+}$, forced deduplication step $S\in\mathbb{Z}^{+}$, temperature decay factor $\alpha\in[0,1]$, pool weight decay factor $\gamma\in[0,1]$
1 $\mathcal{G}=\{G_{i}\}_{i=1}^{k}\leftarrow$ set of subgraphs extracted from $\Pi_{0}$
2 $\Pi\leftarrow\Pi_{0}$, $\Pi^{*}\leftarrow\Pi_{0}$triangleright current and optimal solutions
3 $\mathcal{O}\leftarrow\{\emptyset,\emptyset,\emptyset\}$triangleright set of pools of operators
4 Update-Pools ($\mathcal{G}$, $\mathcal{O}$, $\Pi$) $\triangleright$ initialize the pools
5 $t\leftarrow 1$triangleright temperature in simulated annealing
6 $\mathbf{p}\leftarrow[1,1,1]$
$\triangleright$ pool weights for roulette wheel selection
7 for $i\leftarrow 1,2,...,M$ do
8       $O\sim\mathcal{O}$triangleright sample a pool by the probability $\sigma(\mathbf{p})$
9       update the weight $\mathbf{p}[O]$ of the selected pool $O$ to $(1-\gamma)\cdot\mathbf{p}[O]+\gamma\cdot\max(-\Delta\tau,0)$
10       $\mathbf{h}\leftarrow$ stacked heuristic values for all operators in $O$
11       $o\sim O$triangleright sample an operator by probability $\sigma(\mathbf{h})$
12       $\mathcal{G}^{\prime}\leftarrow$ apply operator $o$ on set $\mathcal{G}$
13       $\Pi^{\prime}\leftarrow$ apply ESTC on each subgraph in $\mathcal{G}^{\prime}$
14       if $\tau(\Pi^{\prime})<\tau(\Pi)$ then
15             update $\mathcal{G}$ to $\mathcal{G}^{\prime}$ and $\Pi$ to $\Pi^{\prime}$
16
17      else
18             update $\mathcal{G}$ to $\mathcal{G}^{\prime}$ and $\Pi$ to $\Pi^{\prime}$ with a probability of $\exp{(\frac{\tau(\Pi)-\tau(\Pi^{\prime})}{t})}$
19
20      call Update-Pools ($\mathcal{G}$, $\mathcal{O}$, $\Pi$, $o$) if $\mathcal{G}$ is updated
21       if $i\,\%\,S=0\;\,\text{or}\;\,\tau(\Pi^{\prime})<\tau(\Pi)$ then
22             call Forced-Deduplication ($\mathcal{G}$, $\mathcal{O}$, $\Pi$)
23
24      update $\Pi^{*}$ to $\Pi$ if $\tau(\Pi)<\tau(\Pi^{*})$
25       update the temperature $t$ to $\alpha\cdot t$
26
27return improved MCPP solution $\Pi^{*}$
28
Algorithm 2 LS-MCPP
Pseudocode (Alg. 2): LS-MCPP begins by computing the corresponding set $\mathcal{G}_{0}$ of subgraphs on the input initial MCPP solution $\Pi_{0}$ [Line 2]. It then initializes a set of pools via a subroutine Update-Pools described in Sec. V-B, each containing dedicated operators of the same type [Lines 2-2]. The pool weight vector $\mathbf{p}$ for the pools and the temperature scalar $t$ are also initialized [Lines 2-2]. The scalar $t$ dynamically adjusts in each iteration [Line 2] to make LS-MCPP progressively harder to accept a non-improving operator over iterations, as per the simulated annealing strategy [van1987simulated] to escape from local minimums. During each iteration [Line 2], vector $\mathbf{p}$ is utilized for roulette wheel selection [goldberg1989genetic] to select the most promising pool $O$ from $\mathcal{O}$ [Line 2], where $\sigma(\mathbf{p})$ denotes the softmax function of $\mathbf{p}$. After selecting pool $O$, its corresponding weight $\mathbf{p}[O]$ is updated with the specified weight decay factor $\gamma$ [Line 2]. Similar to the pool selection, LS-MCPP samples an operator $o$ from the selected pool $O$ with $\sigma(\mathbf{h})$ as the probability function [Line 2]. Here, $\mathbf{h}$ represents the vector of heuristic function $h$ evaluated on all operators in $O$ [Line 2]. Following this, LS-MCPP obtains a new set $\mathcal{G}^{\prime}$ by applying $o$ on the current set $\mathcal{G}$ of subgraphs [Line 2], and then generates the corresponding solution $\Pi^{\prime}$ by applying ESTC on each subgraph in $\mathcal{G}^{\prime}$ [Line 2]. LS-MCPP immediately accepts $o$ with the new $\mathcal{G}^{\prime}$ and the new $\Pi^{\prime}$ if $\Pi^{\prime}$ is better, and otherwise accepts it with a probability of $\exp{(-\Delta\tau/t)}\in[0,1]$ [Lines 2-2]. Upon the acceptance of $o$, LS-MCPP updates the set $\mathcal{O}$ of operator pools via Update-Pools [Line 2]. LS-MCPP also calls a subroutine Forced-Deduplication (described later) every $S$ iteration or when the current solution is better [Line 2]. The iterative search procedure terminates when it reaches the maximum number of iterations $M$, and returns the best solution $\Pi^{*}$ recorded [Line 2].
Forced Deduplication: A critical aspect contributing to the success of LS-MCPP is restricting duplication to only those necessary over the neighborhood exploration, such as those in a narrow path connecting two separate regions of graph $G$. For this purpose, LS-MCPP uses the Forced-Deduplication subroutine as described in Alg. 3 to deduplicate all possible duplication in two folds. The first part iterates through each path $\pi_{i}\in\Pi$ in a cost-decreasing order to remove any U-turns [Lines 3-3]. A U-turn of $\pi_{i}$ is defined as an edge $(u,v)\in E(\pi_{i})$ where $\exists\,(p,q)\in\pi_{i}$ such that $(u,p),(v,q)\in\pi_{i}$, thus forming a “U-turn” pattern of $p\rightarrow u\rightarrow v\rightarrow q$. It follows that we can safely remove $u,v$ to reduce $c(\pi)$ without introducing any new duplication or losing any vertex coverage if $u,v\in V_{i}\cap V^{+}$. Fig. 8 illustrates the procedure to remove U-turns for a path segment progressively. A special case of U-turn often occurs in incomplete hypervertices when $p\rightarrow u\rightarrow v\rightarrow q$ collapses into one edge if $p=v$, where we can safely remove $u,v$ without requiring $v\in V_{i}\cap V^{+}$. The second part [Lines 3-3] recursively applies every deduplicate operator from $O_{d}$ with two prioritizations of (1) higher cost of the corresponding path $\pi_{i}$ and (2) larger heuristic value of the operator. After the above two procedures, it updates all the pools in $\mathcal{O}$ by calling Update-Pools subroutine [Line 3]. It is worth noting that the first part can remove edges that do not constitute deduplicate operators, and some deduplicate operators can remove edges that are not U-turns depending on the MST for ESTC. This complementary relation contributes to the improved performance of LS-MCPP.
Input: set $\mathcal{G}=\{G_{i}\}_{i=1}^{k}$ of subgraphs,      set $\mathcal{O}=\{O_{g},O_{d},O_{e}\}$ of pools,      set $\Pi=\{\pi_{i}\}_{i=1}^{k}$ of coverage paths
1
2for $i\in\operatorname*{arg\,sort}\left(\{c(\pi_{i})\}_{i=1}^{k}\right)$ do
3       while $(u,v)\leftarrow$ any U-turn in $\pi_{i}$ do
4             remove $u,v$ from $G_{i}$ and $\pi_{i}$
5
6
7
8for $i\in\operatorname*{arg\,sort}\left(\{c(\pi_{i})\}_{i=1}^{k}\right)$ do
9       $O\leftarrow\{o_{d}(j,e)\in O_{d}\,|\,i=j\}$
10       for $o\leftarrow$ operator from $O$ with largest heuristic do
11             apply $o$ on $G_{i}$ and update $\pi_{i}$ via ESTC
12
13
14call Update-Pools ($\mathcal{G}$, $\mathcal{O}$, $\Pi$)
Algorithm 3 Forced-Deduplication
Initial Solution: For the choice of the initial solution $\Pi_{0}$ for LS-MCPP, we limit ourselves to existing multi-tree MCPP methods. In addition, the method used to generate $\Pi_{0}$ should also be time-efficient to produce a MCPP solution with good quality. To this end, we choose our initial solution between the solutions resulting from VOR and MFC depending on whose quality is better. Both are integrated with the proposed ESTC algorithm described in Sec. IV-D.
Remark 2.
Our LS-MCPP framework is inherently better suited to utilizing a multi-tree method for generating an initial MCPP solution, which aligns with its design to iteratively improve a set of $k$ subgraphs using a local search strategy and apply ESTC to obtain coverage paths for evaluating the solution cost in each iteration. While a single-tree method can also provide initial subgraphs for LS-MCPP, induced by its solution coverage paths, our empirical finding [tang2024large] indicates that this approach typically results in significant overlaps among subgraphs and tends to reduce the effectiveness of the neighborhood operators within our LS-MCPP framework.
V-B Pools of Operators
We introduce three types of operators: grow operators, deduplicate operators, and exchange operators. These operators are devised to modify the boundaries of each connected subgraph $G_{i}=(V_{i},E_{i})\in\mathcal{G}$ of the input graph $G=(V,E)$. To start with, we define the duplication set $V^{+}=\{v\in V\,|\,n_{v}>1\}$ as the set of vertices covered by more than one path, where $n_{v}=\sum_{i\in I}\left|\{x\in V_{i}|x=v\}\right|$ counts the occurrences of the vertex $v\in V$ in all subgraphs. An operator alters the set $\mathcal{G}$ while ensuring that its property of complete coverage (i.e., $\bigcup_{i\in I}V_{i}=V$) and subgraph connectivity remain invariant. Each operator aims to introduce only necessary duplication, which has the potential to enhance the MCPP solution. Specifically, we consider edge-wise and vertex-wise operations for the above three types of operators. The vertex-wise operator only alters a single vertex one at a time, whereas the edge-wise operator alters a single edge one at a time. In practice, we find that altering an edge $(u,v)\in E$ that connect vertices belonging to different hypervertices (namely $\delta_{u}\neq\delta_{v}$) often brings unnecessary duplications on vertex coverage. Therefore, we restrict the edge-wise operators to only alter edges from a subset of $E$, given as $F=\{(u,v)\in E\,|\,\delta_{u}=\delta_{v}\}\subset E$.
Grow Operator: Let $B_{i}$ denote the set of vertices adjacent to but not part of $V_{i}$, defined as $B_{i}=\{v\in V_{i}\,|\,\exists(u,v)\in E,u\in V\setminus V_{i}\}$. A vertex-wise grow operator $o_{g}^{\text{(v)}}(i,\{v\})$ adds vertex $v\in B_{i}$ and all its adjacent edges to subgraph $G_{i}$. An edge-wise grow operator $o_{g}^{\text{(e)}}(i,\{u,v\})$ expands subgraph $G_{i}$ to add an edge $e=(u,v)\in E_{i}\cap F$ and all its adjacent edges into $G_{i}$, where vertices $u,v\in B_{i}$. We say $o_{g}(i,e)$ is valid iff $\exists(p,q)\in E_{i},(u,p),(v,q)\in E_{i}$. In essence, a valid $o_{g}(i,e)$ can only add an edge $(u,v)$ if there exists a parallel edge $(p,q)$ in $G_{i}$, as shown in the two examples in Fig. 9-(b)(c). This design choice circumvents the introduction of undesirable routing to the coverage path, as illustrated in Fig. 9-(a).
Figure 8: Subroutine Forced-Deduplication executed on a segment of coverage path $\pi_{i}$, where red and black circles represent vertices from $V_{i}\cap V^{+}$ and $V_{i}\setminus V^{+}$, respectively. Frames from left to right: a U-turn $(u,v)\in E(\pi_{i})$ is found, and $u,v$ are removed from $\pi_{i}$ until no U-turn exists.
Figure 9: ESTC path after applying an edge-wise $o_{g}^{\text{(e)}}(i,\{u,v\})$ on subgraph $G_{i}=(V_{i},E_{i})$. Blue circles and black circles represent vertices of the boundary vertex set $B_{i}$ and vertices of $V_{i}$, respectively. (a) An invalid operator without any parallel edge in $G_{i}$. (b)(c) Two valid edge-wise grow operators.
Deduplicate Operator: The deduplicate operator aims to remove unnecessary duplications from a subgraph. For vertex-wise operator $o_{d}^{\text{(v)}}(i,\{v\})$, it removes vertex $v\in V^{+}$ and all its adjacent edges from $G_{i}$. For edge-wise operator $o_{d}^{\text{(e)}}(i,\{u,v\})$, it removes edge $(u,v)\in E_{i}\cap F$ and all its adjacent edges from $G_{i}$, where $u,v\in V_{i}\cap V^{+}$ are already covered by other subgraphs. For edge-wise operator $o_{d}^{\text{(e)}}(i,\{u,v\})$, it is always valid if $\delta_{u}$ contains only $u$ and $v$, as removing them always reduces the cost of the corresponding ESTC path of $G_{i}$ (see Fig. 10-(c)). When $\delta_{u}$ contains not only $u$ and $v$, we say $o_{d}^{\text{(e)}}(i,\{u,v\})$ is valid iff it satisfies the following conditions: (1) any vertex contracted to $\delta_{e}^{t}$ is not in $V_{i}$; (2) all the four vertices contracted to $\delta^{b}_{e}$ are in $V_{i}$; (3) if there exist some $v\in V_{i}$ that is contracted to $\delta_{e}^{l}$ (or $\delta_{e}^{r}$), then all the four vertices contacted to $\delta_{e}^{l}$ (or $\delta_{e}^{r}$) are in $V_{i}$, and all the four vertices contacted to the (only) common neighbor between $\delta_{e}^{l}$ (or $\delta_{e}^{r}$) and $\delta_{e}^{b}$ are in $V_{i}$. The hypervertices $\delta^{t}_{e}$, $\delta^{b}_{e}$, $\delta^{l}_{e}$, and $\delta^{r}_{e}$ denote the four neighbors of $\delta_{u}$ (also equal to $\delta_{v}$) based on the position of $u$ and $v$ within $\delta_{u}$, as shown in Fig. 10-(a) that can be rotated from various symmetric cases. Such validity filtering on the edge-wise deduplicate operators is essential to avoid introducing unnecessary duplications to the ESTC path of the subgraph (see Fig. 10-(b)(c)(d)).
Figure 10: ESTC path after applying an edge-wise $o_{d}^{\text{(e)}}(i,\{u,v\})$ on subgraph $G_{i}=(V_{i},E_{i})$. Red circles and black circles represent vertices in $V_{i}\cap V^{+}$ and not in $V_{i}\setminus V^{+}$, respectively. (a) Four neighbors of the hypervertex $\delta_{u}(=\delta_{v}$). (b) An invalid operator violating the third condition when $\delta_{u}$ is complete. (c)(d) Two valid edge-wise deduplicate operators.
Exchange Operator: Intuitively speaking, an exchange operator combines a grow operator and a deduplicate operator so that the deduplicate operator immediately removes the new duplication introduced by the grow operator, potentially improving the effectiveness of a single operator and speeding up the convergence of the local search. The vertex-wise operator $o_{e}^{\text{(v)}}(i,j,\{v\})$ is constructed from $o_{g}^{\text{(v)}}(i,\{v\})$ and $o_{d}^{\text{(v)}}(j,\{v\})$, where $v\in B_{i}\cap V_{j}$. The edge-wise operator $o_{e}^{\text{(e)}}(i,j,\{u,v\})$ is constructed from $o_{g}^{\text{(e)}}(i,\{u,v\})$ and $o_{d}^{\text{(e)}}(j,\{u,v\})$, Note that the deduplicate operator being combined might not be valid as $v\notin V^{+}$ for vertex-wise ones (or $u,v\notin V^{+}$ for edge-wise ones) are not necessarily covered by other subgraphs before the grow operator is applied. In practice, when updating $O_{e}$ (line 4 of Alg. 4), we iterate each grow operator $o_{g}(i,V_{o_{g}})$ and check if there exists deduplicate operator $o_{d}(j,V_{o_{d}})$ such that $V_{o_{d}}=V_{o_{g}}$ without requiring $V_{o_{d}}\subseteq V^{+}$.
Update Pools: In practice, we find that edge-wise operators work well most of the time but are sometimes unable to be constructed when encountered with many incomplete hypervertices, resulting in early termination of the LS-MCPP search. Vertex-wise operators are less restricted than edge-wise operators, however, tend to introduce unnecessary vertex duplications if applied near complete hypervertices. Therefore, we consider the edge-wise operators primary and make vertex-wise complementary when no edge-wise operators can be found. This is reflected in the Update-Pools function in Alg. 4 to initialize the set $\mathcal{O}$ of pools of operators [Line 4] or update the set $\mathcal{O}$ whenever $\mathcal{G}$ is changed by some operator $o$ [Lines 4-4]. Given the set $\Delta V$ of dirty vertices (defined in line 4 or line 4), Update-Pools first removes all operators involving any $v\in\Delta V$ from each pool [Line 4]. For each $v\in\Delta V$, it then iterates each light subgraph to add a valid edge-wise grow operator to $O_{g}$ if one exists; otherwise, it adds a vertex-wise grow operator [Lines 4-4]. Similarly, deduplicate operators are added to $O_{d}$ by iterating each heavy subgraph [Lines 4-4]. Finally, the pool $O_{e}$ of exchange operators is updated by combining the newly added operators in $O_{g}$ and $O_{d}$ [Line 4].
Input: set $\mathcal{G}=\{G_{i}\}_{i=1}^{k}$ of subgraphs,      set $\mathcal{O}=\{O_{g},O_{d},O_{e}\}$ of pools,      set $\Pi=\{\pi_{i}\}_{i=1}^{k}$ of coverage paths,      [optional] most recent applied operator $o$
1 if $o$ is provided as an input then
2       $V_{o}\leftarrow$ the set of involved vertices in $o$
3       $\Delta V\leftarrow V_{o}\cup\{u\in V\,|\,v\in V_{o},(u,v)\in E\}$
4
5else
6       $\Delta V\leftarrow V$triangleright update for all vertices in $G=(V,E)$
7
8
remove operators from each $O\in\mathcal{O}$ involving any $v\in\Delta V$
9 for $v\in\Delta V$ do
10       for $i\in\{j\in I\,|\,c(\pi_{j})\leq\bar{c}\}$ do
11             if found a valid $o_{g}^{\text{(e)}}=(i,\{u,v\})$ then
12                   $O_{g}\leftarrow O_{g}\cup\{o_{g}^{\text{(e)}}\}$
13            else
14                   $O_{g}\leftarrow O_{g}\cup\{o_{g}^{\text{(v)}}=(i,\{v\})\}$
15
16      for $i\in\{j\in I\,|\,c(\pi_{j})>\bar{c}\}$ do
17             if found a valid $o_{d}^{\text{(e)}}=(i,\{u,v\})$ then
18                   $O_{d}\leftarrow O_{d}\cup\{o_{d}^{\text{(e)}}\}$
19            else
20                   $O_{d}\leftarrow O_{d}\cup\{o_{d}^{\text{(v)}}=(i,\{v\})\}$
21
22
23
update $O_{e}$ by iterating newly added operators in $O_{g},O_{d}$
Algorithm 4 Update-Pools
Properties: After any valid operator is executed, (1) every $G_{i}\in\mathcal{G}$ remains connected since only deduplicate operators can remove edges but a valid deduplicate guarantees that $G_{i}$ remains connected, and (2) the property of the union of the vertices of all subgraphs being the vertex set $V$ is not affected since no vertex coverage is lost. Therefore, with Theorem 3, we know that the coverage paths in an LS-MCPP solution $\Pi$ jointly cover all vertices of $G$ if $\bigcup_{i\in I}V_{i}=V$.
V-C Heuristics for Operator Sampling
To better explore the solution neighborhood, LS-MCPP should carefully determine which operator to sample from the selected pool (line 2 of Alg. 2). To this end, we propose three heuristic functions tailored to the three types of operators to evaluate their potential in guiding the neighborhood search and improving the solution. An operator with a larger heuristic value results in a higher probability of being sampled. For the grow operator, its heuristic function is determined by two considerations: (1) prioritizing growing light subgraphs with smaller path costs using ESTC and (2) prioritizing the edges whose two endpoints have less duplication. Formally, the heuristic value $h(o_{g})$ for a grow operator $o_{g}(i,V_{o_{g}})$ is defined as $h(o_{g})=-k\cdot c(\pi_{i})-\frac{1}{|V_{o_{g}}|}\cdot\sum_{v\in V_{o_{g}}}n_{v}$. Here, $k\geq n_{z}$ holds for any $z\in V$, prioritizing consideration (1) over consideration (2). For deduplicate operator $o_{d}(i,V_{o_{d}})$, the heuristic function is the opposite of the one for grow operators: $h(o_{d})=k\cdot c(\pi_{i})+\frac{1}{|V_{o_{d}}|}\cdot\sum_{v\in V_{o_{d}}}n_{v}$. For exchange operator $o_{e}(i,j,V_{o_{e}})$, its heuristic function is defined as $h(o_{e})=c(\pi_{j})-c(\pi_{i})$ to prioritize two subgraphs with a larger difference between their path costs.
VI Deconflicting MCPP Solutions
In Sec. V, the MCPP solution produced by LS-MCPP does not inherently guarantee that the robots will be conflict-free during their operations in practice. We observe that the robots would be conflict-free in most regions of the graph, as they tend to naturally disperse to cover different regions in an optimized LS-MCPP solution to minimize the overall makespan. Despite this, inter-robot collisions remain a significant challenge in certain congested areas, such as the region of root vertices of the robots when they are clustered or the narrow corridors that only allow one robot to pass through at a time. To tackle this challenge, we leverage techniques from MAPF research, which focuses on finding a set of collision-free paths for a group of robots to navigate across a grid graph. Following the common practice for solving MAPF problems, we develop a two-level hierarchical planner to deconflict an MCPP solution, where the high-level planner determines how the conflicts between the robots are resolved, and the low-level planner focuses on single-robot planning by considering other robots as spatial-temporal obstacles.
VI-A Preliminary
We aim to construct a set $\mathcal{P}=\{P_{i}\}_{i=1}^{k}$ of $k$ conflict-free plans from an input solution $\Pi=\{\pi_{i}\}_{i=1}^{k}$ for the MCPP instance $(G,I,R)$. Each plan $P_{i}=(\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{|P_{i}|})$ is an ordered sequence of states subject to the order of the vertices in $\pi_{i}=(v_{1},v_{2},...,v_{|\pi_{i}|})$. The time-embedded state $\mathbf{x}=(v,t)$ represents the state of a robot arriving vertex $v\in V$ at time $t\geq 0$, and the initial state of each robot $i$ is set as $\mathbf{x}_{1}=(r_{i},0)\in P_{i}$ where $r_{i}\in R$. The time cost $c(P_{i})$ of robot $i$ is the associated time of the last state $\mathbf{x}_{|P_{i}|}$. For convenience, we use $P[j]$, $V(P)\subseteq V$, and $E(P)$ to denote the $j^{\text{th}}$ state $\mathbf{x}_{i}$ of any plan $P$, the set of unique vertices visited by $P$, and the list of (potentially duplicated) edges used by $P$, respectively. For any adjacent vertices $v_{j},v_{j+1}\in\pi_{i}$, a subsequence of plan $P_{i}$ is constructed to transit from state $(v_{j},\cdot)$ to $(v_{j+1},\cdot)$ without any conflict (see Definition 5). To transit from a state $\mathbf{x}=(v,t)$ to another state $\mathbf{x}^{\prime}=(v^{\prime},t^{\prime})$ in a plan, the robot can either wait at vertex $v$ for a time period $t^{\prime}-t>0$ when $v=v^{\prime}$, or move from vertex $v$ to vertex $v^{\prime}\neq v$ via edge $e=(v,v^{\prime})\in E$ at the time cost of $t^{\prime}-t=w_{e}>0$.
Definition 5 (Conflict).
Consider any two plans $P_{a}$ and $P_{b}$ where each state of $P_{a}$ is $\mathbf{x}^{a}_{i}=(v_{i}^{a},t_{i}^{a})$ for $i=1,2,...,|P_{a}|$ and each state of $P_{b}$ is $\mathbf{x}^{b}_{j}=(v_{j}^{b},t_{j}^{b})$ for $j=1,2,...,|P_{b}|$. A conflict on vertex $v\in V$ occurs if there exist $\mathbf{x}^{a}_{i}\in P_{a}$ and $\mathbf{x}^{b}_{j}\in P_{b}$ such that $v^{a}_{i}=v^{b}_{j}=v$ and the two time intervals $[t_{i-1}^{a},t_{i+1}^{a})$ and $[t_{j-1}^{b},t_{j+1}^{b})$ intersects with each other.
Note that two dummy states $\mathbf{x}_{0}=(v_{1},0)$ and $\mathbf{x}_{|P|+1}=(v_{|P|},+\infty)$ are padded before and after each plan $P$ to check if there is a conflict at $\mathbf{x}_{1}$ or $\mathbf{x}_{|P|}$, since the robot associated with $P$ will stay at $v_{1}$ before the task and at $v_{|P|}$ after the task. Our definition of conflict is more conservative than that in most MAPF algorithms. Specifically, we assume that a vertex is blocked not only when it is occupied by a robot, as is typical in the MAPF literature [sharon2015conflict], but also when a robot is moving out of it. By checking for time intervals during which other robots occupy a vertex, we can implicitly identify all conflict types outlined in the literature [stern2019multi], including vertex conflicts, edge conflicts, following conflicts, and swapping conflicts. These conflicts are typically defined for unweighted graphs, where time is discretized into uniform intervals for the robots, allowing their actions to be aligned step by step.
Theorem 7.
It is NP-hard to find an optimal deconflicted MCPP solution with a minimized makespan.
{proof}
The OMG-MAPF problem (see Sec. II-B) was proved to be NP-hard with a makespan minimization objective [mouratidis2024fools], which can be viewed as a special case of our problem that operates solely on unweighted graphs (i.e., $w_{e}=1$ for each $e\in E$). Therefore, it is at least NP-hard to find an optimal deconflicted MCPP solution with a minimized makespan.
Remark 3.
Instead of first solving MCPP and then deconflicting the resulting MCPP solution as proposed, one could also deconflict the MCPP solution at the end of each iteration of LS-MCPP (Alg. 2), which holds the potential to outperform the proposed sequential approach. However, it incurs significantly higher time complexity due to the computational demands of the deconflicting procedure (see Theorem 7 and Tab. II).
VI-B The High-Level Planner
The high-level planner determines how the conflicts defined in Definition 5 are resolved, which builds a search tree by growing child nodes whenever a conflict is detected. At each node of the search tree, a low-level planner is used to compute a plan for each robot with the help of a reservation table to avoid the conflicts that are reserved in it. In short, the reservation table is a dictionary that stores the items with the vertex as key and a set of disjoint time intervals as value.
Figure 11: The high-level Prioritized-Based Search (PBS) planner.
In consideration of time complexity as demonstrated in Theorem 7, we use Priority-based Search (PBS) [ma2019searching] as the high-level planner. PBS maintains a search tree of nodes, where each node $n$ is associated with a set $\mathbf{o}_{n}$ of pairwise partial orders and a set $\mathcal{P}_{n}$ of plans. As illustrated in Fig. 11, PBS resolves the pairwise conflicts between robot $i$ and $j$ at each node $n$ by splitting it into two child nodes, $a$ and $b$, one with the order $i\prec j$ to $\mathbf{o}_{a}$ added and the other with order $j\prec i$ added to $\mathbf{o}_{b}$. For each child node (e.g., node $a$), the corresponding $\mathcal{P}_{a}$ is updated by invoking a low-level planner to update the plan of each robot affected by the addition of the order $j\prec i$ into $\mathbf{o}_{a}$. Each plan is computed with all the plan states from all robots with higher priority reserved in the reservation table. The search tree grows in a depth-first search (DFS) fashion until it reaches a leaf node $n$ where all conflicts in set $\mathcal{P}_{n}$ are resolved using $\mathbf{o}_{n}$. The following remark explains that compared with CBS, PBS is more time efficient due to a much smaller DFS tree, since each search node resolves all the conflicts between the two robots.
Remark 4.
Alternatively, one can employ Conflict-based Search (CBS) [sharon2015conflict] as the high-level planner. CBS addresses pairwise conflicts between robots $i$ and $j$ in each search node by splitting it into two child nodes. In one child node, the plan for robot $i$ is updated by invoking the low-level planner with the conflicted state added to its reservation table; in the other child node, robot $j$’s plan is updated with the conflicted state recorded in its reservation table. As the search tree expands, CBS systematically resolves all possible conflicts. CBS-like algorithms are proven to be optimal for various MAPF variants when the corresponding low-level single-robot planner is optimal. However, this approach often requires longer runtimes to explore a larger number of nodes and construct the search tree necessary to find the optimal solution. This is particularly true in our multi-goal setting of MAPF, where the optimal low-level planner has significantly higher time complexity. According to [mouratidis2024fools], OMG-MAPF, a special case of our deconflicting problem in Sec. VI-A where the edges are unweighted was solved with CBS as the high-level planner, but the method could only scale to sequences of up to 5 goals on relatively small graphs.
VI-C The Low-level Planner
In the lower-level planner, our goal is to compute a partially conflict-free plan $P_{i}$ for robot $i$ to sequentially visit the ordered set of vertices along its path $\pi_{i}\in\Pi$. The plan only avoids conflicts with higher-priority plans by reserving their states in the reservation table. To accommodate planning in graph $G$ with nonuniform edge weights, we adopt the idea from Safe Interval Path Planning (SIPP) [phillips2011sipp] for single-goal shortest path planning. In essence, SIPP expands upon the heuristic search concept of space-time A [hart1968formal] with the only difference on how the states are stored in the reservation table. Given a state $\mathbf{x}=(v,t)$ to be recorded in the reservation table, space-time A records the discrete time $t$ for vertex $v$, whereas SIPP records the time interval $[t^{\prime},t)$ for vertex $v$ where $t^{\prime}$ is the arrival time of the preceding state $\mathbf{x}^{\prime}=(v^{\prime},t^{\prime})$ of $\mathbf{x}$. Inspired by [mouratidis2024fools], we develop three approaches for the low-level planner upon SIPP, namely chaining approach, holistic approach, and a novel adaptive approach. Before the low-level planning, we do a preprocessing step on each $\pi_{i}\in\Pi$ to remove any $v\in R\setminus\{r_{i}\}$ from $\pi_{i}$. This is crucial to make the low-level planner more robust and avoid a doomed failure when a root vertex of another robot is an intermediate goal in the goal sequence.
Figure 12: A PBS node with partial order $2\prec 1$. Left: first conflict found at vertex $z$ between state $(z,3)$ of robot $1$ and state $(z,2)$ of robot $2$, where the edge weights are annotated along the edge, and the two paths $\pi_{1}$ and $\pi_{2}$ are colored in red and blue, respectively. Right: plans $P_{1}$ and $P_{2}$ of the two robots and the reservation table $\text{rt}_{1}$ of robot $1$. The chaining approach cannot resolve the conflict using SIPP, as both $u$ and $z$ are unreachable once robot $1$ reaches state $\mathbf{x}=(v,2)$. The holistic and adaptive approaches can resolve the conflict using ML-SIPP when reaching state $\mathbf{x}$ by backtracking to states expanded from $(u,0)$.
Chaining Approach: Given a path $\pi=(v_{1},v_{2},...,v_{|\pi|})$ for any robot, the chaining approach considers each pair of consecutive vertices of $\pi$ as a single-robot path planning problem from a goal state $\mathbf{x}_{i}^{(\pi)}=(v_{i},\cdot)$ to the next goal state $\mathbf{x}_{i+1}^{(\pi)}=(v_{i+1},\cdot)$. We then use SIPP with the admissible true distance heuristic [silver2005cooperative], denoted as $h^{T}(\mathbf{x},\mathbf{x}^{(\pi)})$. It requires an auxiliary backward search from the goal state $\mathbf{x}^{(\pi)}$ to compute the minimum cost to any state $\mathbf{x}$ ignoring any dynamic obstacles in the time dimension. Note that reaching a goal state $\mathbf{x}^{(\pi)}$ in SIPP is determined by the vertex regardless of its arrival time. By concatenating the resulting consecutive plans one after the other, we obtain the final plan $P_{i}$ for the robot $i$. According to [mouratidis2024fools], the chaining approach can be shown to be suboptimal and also not complete. Such property persists when we directly replace the single-robot planner of space-time A with SIPP, as demonstrated in the example of Fig. 12. Intuitively speaking, its suboptimality and incompleteness result from its restricted search space.
Holistic Approach: The holistic approach borrows the similar idea from Multi-Label A [grenouilleau2019multi, zhong2022optimal], originally for multi-goal MAPF variants in unweighted graphs. By incorporating a label dimension into the robot state, we adopt Multi-Label A for our weighted graph settings with SIPP, namely Multi-Label Safe-Interval Path Planning (ML-SIPP). Specifically speaking, for each vertex $v_{i}$ in the path $\pi=(v_{1},v_{2},...,v_{|\pi|})$, we define its corresponding goal state as $\mathbf{x}_{i}^{(\pi)}=(v_{i},\cdot,l_{i})$ with label $l_{i}=i$ indicating $v_{i}$ as its current goal vertex to reach. Here, the arrival time can be arbitrary, consistent with the goal-reaching condition on only the vertex dimension as in the chaining approach. The transition between any two consecutive states $\mathbf{x}_{j}$ and $\mathbf{x}_{j+1}$ in a plan is thereby updated for the label dimension, regardless of the inducing actions: if $\mathbf{x}_{j}$ is a goal state with label $l$, the label of $\mathbf{x}_{j+1}$ is set as $l+1$; otherwise, the label inherits the label $l$ of $\mathbf{x}_{j}$. To further tighten the true distance heuristics $h^{T}$ in SIPP, the same concatenated heuristics from Multi-Label A for ML-SIPP:
$h(\mathbf{x},\mathbf{x}_{i})=h^{T}(\mathbf{x},\mathbf{x}_{i}^{(\pi)})+\sum_{j=i}^{|\pi|-1}h^{T}(\mathbf{x}_{j}^{(\pi)},\mathbf{x}_{j+1}^{(\pi)}),$ (5)
where $\mathbf{x}_{i}^{(\pi)}$ is the current goal state to reach updated throughout the search procedure. One can easily verify that $h$ is admissible as $h^{T}$ is admissible, with which we can prove the holistic approach is an optimal and complete low-level planner.
Adaptive Approach: So far we have introduced two approaches of the low-level planner: the chaining approach is fast yet incomplete, whereas the holistic approach is complete and optimal yet has large time and space complexity with the additional label dimension. To improve the practicability of both approaches, we propose a novel adaptive low-level planner in Alg. 5 combining their benefits. Instead of running ML-SIPP on the entire sequence of goals as in the holistic approach, our adaptive approach calls ML-SIPP on a limited window of the goal sequences that traces back to previous goals whenever a sequence containing only the current goal is unsolvable by ML-SIPP [Lines 5-5]. The adaptive windows size is restricted by the input max backtracking size parameter $b_{\text{max}}\in\mathbb{Z}^{+}$. When ML-SIPP finds a valid plan for the sequence containing only the current goal, it proceeds to the next goal and resets the current backtracking window $b$ [Lines 5-5]. Note that ML-SIPP falls back to SIPP when the sequence $S_{\text{goal}}$ of goals states contains only one state. Our adaptive approach is complete since it returns the ML-SIPP solution to sequentially visit the goals in $\pi$ in the worst-case scenario [Lines 5-5]. In practice, we find that setting $b_{\text{max}}$ to a small integer is sufficient enough to resolve all conflicts even with a limit on the number of expanded nodes, as we will see in Sec. VII.
Input: path $\pi=(v_{1},v_{2},...,v_{|\pi|})$ for robot $i$
Param: max backtracking size $b_{\text{max}}\in\mathbb{Z}^{+}$
1 remove every $v\in R\setminus\{r_{i}\}$ from $\pi$
2 $\mathbf{x}_{\text{last}}\leftarrow\left(v_{1},0\right)$,  $j=2$,  $b\leftarrow 1$
3 $P\leftarrow\left(\mathbf{x}_{\text{last}},\right)$,  $S_{\text{goal}}\leftarrow(\mathbf{x}^{(\pi)}_{j}=(v_{j},\cdot),\,)$
4 while $j\leq|\pi|$ do
5       find a plan from $\mathbf{x}_{\text{last}}$ to $S_{\text{goal}}$ via ML-SIPP
6       if a valid plan $\tilde{P}$ is found then
7             append $\tilde{P}$ to the end of $P$
8             $j\leftarrow j+1$,  $b\leftarrow 1$
9             $S_{\text{goal}}\leftarrow(\mathbf{x}^{(\pi)}_{j}=(v_{j},\cdot),\,)$
10
11       else if $b\leq b_{\text{max}}$ or $P$ is not empty then
12             pop the last state from $P$ until $\mathbf{x}_{j-b}^{(\pi)}$
13             $S_{\text{goal}}\leftarrow(\mathbf{x}_{j-b}^{(\pi)},\,)+S_{\text{goal}}$,  $b\leftarrow b+1$
14
15       else
16             $\mathbf{x}_{\text{last}}\leftarrow(v_{1},0)$, $S_{\text{goal}}\leftarrow(\mathbf{x}_{1}^{(\pi)},...,\mathbf{x}_{|\pi|}^{(\pi)})$
17             return plan from $\mathbf{x}_{\text{last}}$ to $S_{\text{goal}}$ via ML-SIPP
18
19
20      $\mathbf{x}_{\text{last}}\leftarrow$ the last state in $P$
21
return $P$
Algorithm 5 Adaptive-Low-Level-Planner
VII Experiments and Results
This section describes our numerical results of simulations on an Ubuntu 22.04 PC with a 2.50GHz Intel® Core® i5-13490F CPU and 32GB RAM. Our proposed algorithms and the baseline methods are publicly available at Github [^1] https://github.com/reso1/LS-MCPP, with all the benchmark instances and reported numerical results provided as well. In Sec. VII-E, we deployed the proposed algorithms on two Pepper [^2] https://www.aldebaran.com/en/pepper robots in an indoor environment to validate our planning pipeline for MCPP.
VII-A Experiment Setup
Instances: We generate a set of instances to benchmark different methods utilizing the four-way connected 2d undirected grid graphs from previous grid-based MCPP works [tang2023mixed, tang2024large] as well as the 2D pathfinding benchmark [sturtevant2012benchmarks]. As shown in Fig. 13, the graphs in row 1 have unweighted edges with a uniform weight of 1, whereas the graphs in row 5 have weighted edges whose weights are randomly generated from 1 to 3. Originally, each graph can be contracted to a hypergraph where all its hypervertices are complete. The original graphs are then used to generate 12 groups of 12 mutated MCPP instances (i.e., mutations) of $k$ robots, resulting from a removal ratio $\rho=0,1,...,11$ and a seed $s=0,1,...,11$. Each mutation $(G,I,R)$ with $\rho$ and $s$ is constructed by (1) randomly sampling (with seed $s$) a set of $k$ vertices from $G$ as $R$ and (2) randomly removing (with seed $s$) vertices from $G$ by $x=(\frac{100}{12}\cdot\rho)\%$, meanwhile ensuring the root vertex of any robot and all its neighbors are not removed. The root vertices in $R$ are sampled to ensure they are not cut vertices for constructing well-formed instances [vcap2015complete, ma2019searching], that is, a robot can wait indefinitely at its root vertex without blocking any other robot. Note that the graph edge weights of each mutation are also randomly generated using its corresponding seed $s$. For all mutations, we set $C=0.5$ as the $90^{\circ}$ turning cost as described in Sec. LABEL:sec:turn_cost.
Baselines: We describe all the methods being compared in our experiments. (1) Single-Robot CPP: To verify the effectiveness of our ESTC with the proposed hyperedge weight definition, we use a baseline method, namely ESTC-UW, that computes the circumnavigating coverage path on a spanning tree assuming the edges are unweighted. We also performed an ablation study on the two local optimizations for ESTC: parallel rewiring (PR) in Sec. IV-A and turn reduction (TR) in Sec. IV-C. (2) MCPP: As mentioned in Sec. IV-D, the proposed ESTC algorithm can be integrated into various grid-based MCPP methods, making them immediately applicable to MCPP. To reiterate, VOR, MFC, MIP, and MSTC are the comparing baseline methods with our LS-MCPP. In addition, we also compare LS-MCPP with and without the vertex-wise operators that are new to our previous work [tang2024large]  [^3] In the ablation study of our previous work [tang2024large], we have shown the effectiveness of grow, deduplicate, and exchange operators as well as the use of Forced Deduplication (Alg. 3) and heuristic sampling for LS-MCPP., namely LS(+VO) and LS(-VO) respectively. Note that all the methods compared use ESTC with TR and PR applied by default. (3) Deconflicted MCPP: Given LS(+VO) solutions to produce the initial time-embedded trajectories, we compare the deconflicting performance with the chaining approach (CHA), the multi-label approach (MLA), and the adaptive approach (ADA) as the low-level planner and PBS as the high-level planner.
Parameters: Given an MCPP instance $(G,I,R)$, we describe the settings of all the parameters used in the simulation experiments. For our LS-MCPP (Alg. 2), we set the maximum number of iterations $M=1.0e^{3}\times\sqrt{|V|/k}$ for a good tradeoff between planning efficiency and solution quality, the forced deduplication step size $S=\lfloor M/20\rfloor$ to arrange at least 20 Forced-Deduplication calls, the temperature decay factor $\alpha=\exp{\frac{\log 0.2}{M}}$ to decrease the temperature $t$ from $1$ to $0.2$, and the pool weight decay factor $\gamma=1.0e^{-2}$. For MIP, we set a runtime limit of one hour for the mixed-integer program solving. For the Deconflicted MCPP algorithms in Sec. VI-C, we set a total runtime limit of one hour for the overall procedure. Specifically for the adaptive approach (Alg. 5), we find that setting the max backtracking size $b_{\text{max}}=5$ is sufficient to resolve all local conflicts efficiently.
VII-B Ablation Study
Figure 13: Ablation Study on 6 unweighted graphs (row 1) and 6 weighted graphs (row 5). Each plot (rows 2-4 and 6-8) includes 12 groups of mutated instances with $k$ robots, each contains 12 mutations that randomly remove a portion (x-axis) of vertices, with random root vertices and edge weights (see Sec. VII-A). Single-Robot CPP (rows 2 and 6): mean and standard deviation (std) of cost reductions of ESTC variants over ESTC-UW. MCPP (rows 3 and 7): mean and std of makespan reductions of MCPP methods over VOR. Deconflicted MCPP (rows 4 and 8): mean and std of makespan increments over the input LS-MCPP solution for mutations where all approaches have at least one mutation succeeded (left y-axis) and their success rates in all 12 mutations (right y-axis). Note that when one or two approaches fail at all 12 mutations, we also plot for the mutations where the other approach(es) all succeed.
Single-Robot CPP: Comparing all four ESTC variants with the baseline ESTC-UW, we observe a more critical cost reduction on weighted graphs (row 6 of Fig. 13) than on unweighted graphs (row 2 of Fig. 13). For unweighted graphs, TR results in significant performance improvements since uniformly weighted edges can be aligned in one direction to minimize turning costs. For weighted graphs, TR is much less effective as there are few tie-breaking cases for reducing the turning costs, while PR achieves higher cost reductions as the vertex removal percentage increases. In general, ESTC shows its effectiveness for single-robot CPP, and both PR and TR contribute to substantial path cost reductions.
Figure 14: LS-MCPP solutions with increasing max iterations $M=s\times\sqrt{|V|/k}$. Colored lines represent different values of $s$.
Table I: Runtime Comparison of MCPP Methods.
floor
small
floor
medium
floor
large
ht_ch
antry
AR02
05SR
Shang
hai2
VOR 0.01$\pm$0.01s 0.03$\pm$0.01s 0.1$\pm$0.01s 0.2$\pm$0.03s 0.4$\pm$0.04s 1.5$\pm$0.1s
MFC 0.02$\pm$0.01s 0.21$\pm$0.02s 1.0$\pm$0.08s 6.8$\pm$0.11s 20$\pm$1.35s 4.7$\pm$0.3m
MIP 1h 1h / / / /
MSTC* 0.01$\pm$0.01s 0.12$\pm$0.05s 0.6$\pm$0.24s 3.5$\pm$1.63s 9.6$\pm$4.4s 1.7$\pm$0.7m
LS(-VO) 2.38$\pm$2.50s 0.57$\pm$0.2m 0.9$\pm$0.1m 1.8$\pm$0.2m 3.0$\pm$0.3m 9.9$\pm$1.5m
LS(+VO) 2.87$\pm$1.89s 0.69$\pm$0.1m 0.9$\pm$0.1m 2.1$\pm$0.2m 3.4$\pm$0.3m 13$\pm$0.8m
terrain
small
terrain
medium
terrain
large
ost002d
AR07
01SR
New
York1
VOR 0.01$\pm$0.01s 0.04$\pm$0.01s 0.1$\pm$0.01s 0.4$\pm$0.04s 0.6$\pm$0.1s 1.6$\pm$0.1s
MFC 0.04$\pm$0.01s 0.40$\pm$0.04s 1.1$\pm$0.22s 15$\pm$1.46s 0.7$\pm$0.1m 5.4$\pm$0.2m
MIP 1h 1h / / / /
MSTC* 0.04$\pm$0.02s 0.50$\pm$0.21s 1.4$\pm$0.55s 15$\pm$6.71s 0.5$\pm$0.3m 2.6$\pm$0.8m
LS(-VO) 5.88$\pm$3.41s 0.46$\pm$0.1m 0.6$\pm$0.2m 2.6$\pm$0.3m 4.9$\pm$0.5m 9.1$\pm$0.5m
LS(+VO) 6.65$\pm$2.86s 0.51$\pm$0.1m 0.7$\pm$0.1m 3.0$\pm$0.3m 5.7$\pm$0.4m 13$\pm$0.5m
MCPP: We compare the solution quality of MCPP between different methods in rows 3 and 7 of Fig. 13. Generally speaking, the mutated instances become more complex as the vertex removal ratio increases. Our LS-MCPP consistently outperforms other methods across all instances, while the performance of the other methods varies, with each showing advantages in different instances. Note that due to the huge memory and runtime requirements of MIP, we only test it on relatively smaller instances (i.e., names with a suffix of small or medium), and its performance is highly restricted by the runtime limit of 1 hour. Specifically for our LS-MCPP, as the vertex removal ratio increases, LS(+VO) performs better with the additional vertex operators compared to LS(-VO). However, when the instances become more complex with a large map and more robots, the solution quality of LS(+VO) gets closer to LS(-VO) due to an inadequate number of local search iterations limited by $M$. To demonstrate that we can further improve the solution quality by saturating LS(+VO) with more iterations, we compare the LS(+VO) solutions with increasing maximum numbers of iterations, as shown in Fig. 14. The solution quality steadily improves as the maximum number $M$ of iterations increases with a linear increase in runtime. In practice, $M$ can be flexibly adjusted according to the available runtime budget, allowing a trade-off between better solution quality and faster planning.
Figure 15: Deconflicting different MCPP solutions using PBS as the high-level planner and our adaptive approach as the low-level planner.
Deconflicted MCPP: We explain the Deconflicted MCPP ablation study results in Fig. 13 (rows 4 and 8). CHA demonstrates the highest makespan increments and the lowest success rate due to its inability to backtrack to previous goal states, resulting in incompleteness. This limitation makes CHA susceptible to ”dead-ends” when no valid succeeding states are available. Consequently, it reports a higher number of failure nodes during PBS search, which either prevents PBS from finding a valid priority ordering or significantly increases the number of PBS nodes explored (sometimes exceeding the runtime limit in large-scale instances). MLA produces high-quality trajectory sets, with no makespan increases for most of the ”easier” instances it can handle within the runtime limit. However, for more complex instances involving larger maps and more robots, the added label dimension across the entire coverage path expands the search space considerably, which significantly increases planning time for each PBS node and results in a lower success rate due to timeouts. Our ADA succeeds in all instances, achieving high solution quality with no makespan increases in most cases. This is due to its ability to restrict the label-augmented search space within a local window, rather than across the entire coverage path as MLA, and only when backtracking to previous states is necessary. In Fig. 15, we also compare the deconflicting results of different MCPP solutions for floor_large and terrain_large using PBS and ADA, following the same plotting conventions as in rows 4 and 8 of Fig. 13. The results demonstrate the better solution quality of the proposed LS-MCPP when used as the input MCPP solution for deconfliction. In contrast, as MFC and MSTC* have larger makespans due to the higher overlapping covered regions across the map, the deconflicting procedures on them are more time-consuming. Notably, there is one failure case of deconflicting the MFC solution where the one-hour runtime limit is exceeded. For VOR, as there are no overlapping regions in its solution, PBS immediately returns the initial time-embedded trajectories without any conflict.
VII-C Planning Runtime Comparison
We report the planning runtime of all the MCPP methods and the Deconflicted MCPP methods described in Sec. VII-A across all the 12 groups of instances mutated from the grid graphs shown in rows 1 and 5 of Fig. 13. We show the planning runtime comparison of methods for MCPP in Tab. I, where our LS-MCPP requires more time to explore the solution neighborhood. In particular, adding vertex-wise operators only increases the runtime of LS-MCPP by a little if we compare LS(-VO) and LS(+VO). Nevertheless, it produces high-quality MCPP solutions within around 13 minutes even for large-scale instances such as NewYork1 and Shanghai2, which is acceptable for offline planning in real-world robotics applications.
In Tab. II, we compare the planning runtime of the three low-level planners with PBS as the high-level planner, using the LS(+VO) solution as input. Each approach-instance data cell displays the runtime for instances where all three approaches succeed at the top, and the runtime for instances where each approach succeeds is shown at the bottom. Note that for the top item, when one or more approaches fail for all 12 instances, we also include the runtime for instances where the other approaches all succeed. We can see that for the relatively easier instances where all three approaches succeed at least one mutation, CHA is slightly faster than ADA, whereas MLA is the slowest. For more complex instances (e.g., NewYork1) where CHA only succeeds at a few mutations (see Fig. 13), ADA can still produce a set of valid trajectories for all of them within 10 minutes, and sometimes even faster than CHA who needs to explore more PBS nodes to find a successful set of trajectories.
Table II: Runtime Comparison of Deconflicted MCPP Methods.
floor
small
floor
medium
floor
large
ht_ch
antry
AR02
05SR
Shang
hai2
CHA 0.10$\pm$0.1s 0.7$\pm$0.2s 6.9$\pm$2.7s 0.5$\pm$0.7m 1.6$\pm$1.5m 9.0$\pm$8.4m
0.10$\pm$0.1s 0.7$\pm$0.2s 6.9$\pm$2.7s 0.9$\pm$2.7m 1.5$\pm$1.4m 9.0$\pm$8.4m
MLA 0.16$\pm$0.1s 4.8$\pm$3.9s 3.7$\pm$4.4m 25$\pm$23m 3.2$\pm$0.0m /
0.17$\pm$0.2s 5.2$\pm$4.5s 4.0$\pm$4.8m 34$\pm$21m 31$\pm$28m /
ADA 0.10$\pm$0.1s 0.7$\pm$0.2s 7.3$\pm$3.2s 0.4$\pm$0.5m 1.4$\pm$1.4m 9.1$\pm$6.3m
0.10$\pm$0.1s 0.8$\pm$0.2s 8.1$\pm$4.1s 0.6$\pm$0.6m 1.4$\pm$1.2m 7.0$\pm$5.2m
terrain
small
terrain
medium
terrain
large
ost002d
AR07
01SR
New
York1
CHA 0.18$\pm$0.1s 1.3$\pm$0.6s 9.3$\pm$3.7s 1.0$\pm$0.6m 2.4$\pm$1.6m 5.8$\pm$7.0m
0.18$\pm$0.1s 1.3$\pm$0.6s 9.4$\pm$3.8s 1.3$\pm$1.1m 2.4$\pm$1.6m 5.8$\pm$7.0m
MLA 0.27$\pm$0.2s 5.9$\pm$13s 5.4$\pm$12m 24$\pm$24m / /
0.28$\pm$0.3s 5.8$\pm$12s 5.2$\pm$12m 38$\pm$21m / /
ADA 0.20$\pm$0.1s 1.3$\pm$0.5s 9.9$\pm$4.7s 0.9$\pm$0.5m 2.1$\pm$1.0m 4.4$\pm$3.0m
0.20$\pm$0.2s 1.3$\pm$0.5s 10$\pm$4.6s 1.2$\pm$0.6m 2.1$\pm$0.8m 4.2$\pm$2.4m
VII-D Physical Robot Experiments
Figure 16: Demonstration of Cayenne and Cumin performing the coverage task cooperatively without conflict. Four frames from the task execution are illustrated with their trajectories and the graph shown on the left side. Static obstacles, trajectory of Cayenne, and trajectory of Cumin are highlighted in red, blue, and yellow respectively. The entire planning pipeline was completed in 0.1 seconds and the coverage task was completed in approximately 6 minutes.
We constructed an MCPP instance of an indoor environment with a Vicon [^4] https://www.vicon.com motion capture system for localization as shown in Fig. 17. Our pipeline first computes the conflict-free set of trajectories for the robots by first finding an MCPP solution via LS-MCPP and then deconflicting the MCPP solution via PBS and ADA. After offline planning is completed, the trajectories are sent to the robots to perform the coverage task through a centralized scheduler written in ROS [^5] https://www.ros.org. Fig. 17 demonstrates the execution procedure of the resulting trajectories using our planning pipeline. The detailed setup is described as follows.
Multi-Robot System: We employ two Pepper robots, named Cayenne and Cumin, each with a square bounding box of 0.55 meters. The 2D grid graph represents the indoor environment as a $10\times 6$ map with each grid unit sized at 0.65 meters to accommodate their bounding boxes. The initial root vertices of Cayenne and Cumin are placed at the bottom-left and top-right grids of the graph, respectively. Both Peppers can have omnidirectional movement driven by velocity commands $u=(v_{x},v_{y},\omega)$. The linear velocities $v_{x}$ and $v_{y}$ are capped at 0.15 meters per second, while the angular velocity $\omega$ is restricted to 0.5 radians per second. To determine Pepper’s non-uniform traversal costs in the graph (i.e., the edge weights), we recorded the maximum and minimum time costs to traverse a single grid unit by performing random move actions across the graph. We then sampled weights randomly within this range and assigned them to the edges. Although Peppers can move omnidirectionally, we assume they always face the moving directions by integrating turn actions as described in Sec. LABEL:sec:turn_cost. We set $C$ as the maximum action time of 90 rotations for Peppers. In the final constructed graph, the time costs for move actions range from 7.2 to 10.8 seconds, while the time cost for a 90 turn is 6.0 seconds.
Centralized Scheduler: Given the set of conflict-free trajectories resulted from our planning pipeline, we generate an action sequence consisting of move, turn, wait actions between consecutive states for each Pepper. A centralized scheduler is employed to sequentially send actions to the corresponding Pepper, maintaining a positional tolerance of 0.01 meters and a rotational tolerance of 0.02 radians. The velocity commands for each action are generated using a proportional controller operating at 20 Hz to align with the nominal action time cost. In practice, the move and turn action time costs are not precise for the Peppers, as we assume instantaneous velocity changes and do not account for their real-time dynamics or potential perturbations. This often leads to an overestimation of the actual action time costs, as the time for acceleration and deceleration is ignored, though underestimations can also occur occasionally. Therefore, we make the centralized scheduler to synchronize the discrepancies between the nominal time costs $t$ and the actual time costs $t^{\prime}$ for the actions. Upon completing each action, Pepper sends the time difference $\Delta t=t-t^{\prime}$ back to the scheduler. If $\Delta t<0$, the scheduler signals Pepper to wait for a period of $-\Delta t$ to match the nominal time cost; otherwise, the scheduler instructs the other Pepper to add $\Delta t$ to its waiting period after completing its current action.
Remark 5.
Our centralized scheduler currently uses a basic approach to synchronize time differences between nominal and actual action time costs, which may lead to increased coverage task completion times. More advanced techniques from MAPF execution research, such as constructing an action dependency graph [honig2019persistent], could address this issue. However, exploring these advanced methods is beyond the scope of this paper.
VII-E Physical Robot Experiments
Figure 17: Demonstration of Cayenne and Cumin performing the coverage task cooperatively without conflict. Four frames from the task execution are illustrated with their trajectories and the graph shown on the left side. Static obstacles, trajectory of Cayenne, and trajectory of Cumin are highlighted in red, blue, and yellow respectively. The entire planning pipeline was completed in 0.1 seconds and the coverage task was completed in approximately 6 minutes.
We deployed the proposed algorithmic pipeline for MCPP on two Pepper [^6] https://www.aldebaran.com/en/pepper robots in an indoor environment with a Vicon [^7] https://www.vicon.com/ motion capture system for localization. Given the grid graph of the indoor environment, the pipeline computes the conflict-free set of plans for the robots by first finding an MCPP solution via LS-MCPP and then deconflicting the MCPP solution via PBS and ADA. After offline planning is completed, the plans are sent to the robots to perform the coverage task through a centralized scheduler written in ROS [^8] https://www.ros.org/. Fig. 17 demonstrates the execution procedure of the resulting plans using our planning pipeline, and the detailed experiment setup is described as follows.
Multi-Robot System: We employ two Pepper robots, named Cayenne and Cumin, each with a square bounding box of 0.55 meters. The 2D grid graph represents the indoor environment as a $10\times 6$ map with each grid unit sized at 0.65 meters to accommodate their bounding boxes. The initial root vertices of Cayenne and Cumin are placed at the bottom-left and top-right grids of the graph, respectively. Both Peppers are capable of omnidirectional movement (i.e., holonomic) driven by velocity commands $u=(v_{x},v_{y},\omega)$. The linear velocities $v_{x}$ and $v_{y}$ are capped at 0.15 meters per second, while the angular velocity $\omega$ is restricted to 0.5 radians per second. To determine Pepper’s non-uniform move action costs in the graph (i.e., the edge weights $w_{e}$), we recorded the maximum and minimum time costs to traverse a single grid unit by performing random move actions across the graph. We then sampled weights randomly within this range and assigned them to the edges. We set $C$ as the maximum time required for 90 rotations. In the final constructed graph, the time costs for move actions range from 7.2 to 10.8 seconds, while the time cost for a 90 turn is 6.0 seconds.
Centralized Scheduler: Given the set $\mathcal{P}$ of conflict-free plans resulted from our planning pipeline, we generate an action sequence consisting of move, turn, wait actions between consecutive states for each Pepper. A centralized scheduler is employed to sequentially send actions to the corresponding Pepper, maintaining a positional tolerance of 0.01 meters and a rotational tolerance of 0.02 radians. The velocity commands for each action are generated using a proportional controller operating at 20 Hz to align with the nominal action time cost. In practice, the move and turn action time costs are not precise for the Peppers, as we assume instantaneous velocity changes and do not account for their real-time dynamics or potential perturbations. This often leads to an overestimation of the actual action time costs, as the time for acceleration and deceleration is ignored, though underestimations can also occur occasionally. Therefore, we make the centralized scheduler to synchronize the discrepancies between the nominal time costs $t$ and the actual time costs $t^{\prime}$ for the actions. Upon completing each action, Pepper sends the time difference $\Delta t=t-t^{\prime}$ back to the scheduler. If $\Delta t<0$, the scheduler signals Pepper to wait for a period of $-\Delta t$ to match the nominal time cost; otherwise, the scheduler instructs the other Pepper to add $\Delta t$ to its waiting period after completing its current action.
Remark 6.
Our centralized scheduler currently uses a basic approach to synchronize time differences between nominal and actual action time costs, which may lead to increased coverage task completion times. More advanced techniques from MAPF execution research, such as constructing an action dependency graph, could address this issue. However, exploring these advanced methods is beyond the scope of this paper.
VIII Conclusions & Future Work
In this paper, we presented a comprehensive approach to MCPP on 2D grid graphs, addressing key challenges including grids with incomplete quadrant coarsening, non-uniform traversal costs, inter-robot conflicts, and turning costs in large-scale, real-world scenarios. Our contributions began with the ESTC paradigm, designed to overcome limitations of standard STC-based paradigms by enabling path generation on any grid structure, even those with incomplete quadrant coarsening. We further enhanced ESTC with two local optimizations, improving both path quality and computational efficiency. Building on ESTC, we developed LS-MCPP, a local search framework that integrates ESTC to iteratively refine and optimize coverage paths, yielding high-quality solutions for MCPP on arbitrary grid graphs. To address inter-robot conflicts, we introduced a new MAPF variant as a post-processing step for MCPP. This approach marks the first integration of MAPF techniques into MCPP, bridging two critical fields in multi-robot coordination. Our solver effectively resolves conflicts while preserving solution quality, demonstrating robust performance even in complex environments with up to 100 robots. Our experimental results confirm that the proposed pipeline is scalable, robust, and capable of delivering high-quality solutions, making it suitable for a wide range of real-world multi-robot applications across diverse environments. This work lays a strong foundation for advancing multi-robot systems with implications for practical deployments in real-world applications. Future work will explore encoding action costs on grid graphs with more precise action dynamics, further accelerating the local search through parallelization and machine learning techniques, and integrating online replanning to handle real-time changes and execution deviations.